### v1

# OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models

Abstract
Spatial reasoning is a key aspect of cognitive psychology and remains a major bottleneck for current vision-language models (VLMs). While extensive research has aimed to evaluate or improve VLMs’ understanding of basic spatial relations, such as distinguishing left from right, near from far, and object counting, these tasks represent only the most fundamental level of spatial reasoning. In this work, we introduce OmniSpatial, a comprehensive and challenging benchmark for spatial reasoning, grounded in cognitive psychology. OmniSpatial covers four major categories: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, with 50 fine-grained subcategories. Through Internet data crawling and careful manual annotation, we construct over 1.5K question-answer pairs. Extensive experiments show that both open- and closed-source VLMs, as well as existing reasoning and spatial understanding models, exhibit significant limitations in comprehensive spatial understanding. We further analyze failure cases and propose potential directions for future research.
Overview of OmniSpatial Benchmark.
1 Introduction
Spatial reasoning plays a crucial role in bridging visual observation to robotic action [43, 88, 89], autonomous driving, and AR/VR. For models to execute tasks effectively, they must understand spatial relationships to determine appropriate actions. To enhance the spatial understanding of Vision-Language Models (VLMs), prior works [15, 17, 11, 68, 102, 123, 121, 89] have integrated spatial information into datasets, enabling basic forms of spatial reasoning. Various benchmarks [26, 104, 101, 48, 34, 102, 121, 89] have been introduced to systematically evaluate such capabilities, focusing on tasks like recognizing left and right, estimating depth, and constructing cognitive maps [121, 106, 72]. Additionally, spatial reasoning has been applied to manipulation tasks [25, 8, 131, 20, 123], allowing systems to position objects according to specified spatial rules [89].
Despite these advances, existing approaches largely focus on basic spatial understanding, which typically includes identifying simple spatial relations such as relative positions (left, right, front, back), proximity (near, far), and object-level counting. Recent models, including state-of-the-art reasoning systems like ChatGPT o3 [79] and Gemini-2.5-Pro [97], have achieved over 90% accuracy on benchmarks such as SpatialBot-Bench [11] and EmbSpatial [26], suggesting that these basic tasks are approaching saturation, as shown in OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models.
However, complex spatial reasoning remains a significant challenge [35, 5, 83, 50, 82, 16]. Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships [7, 109, 95]. For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route. Similarly, tasks like inserting a knife into a rack or flattening a box demand reasoning about object rotation, deformation, and spatial compatibility—far beyond static object placement.
From the perspective of cognitive psychology, complex spatial reasoning extends beyond basic relations and includes dynamic world knowledge reasoning, interactive spatial behavior with environments or agents, logical analysis of 3D spatial structures, and perspective-taking abilities.
Therefore, to address these limitations and systematically assess higher-order spatial reasoning, we propose OmniSpatial, a comprehensive benchmark capturing the breadth and depth of spatial cognition. OmniSpatial categorizes spatial reasoning into four key dimensions: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, providing a foundation for advancing the next generation of spatially-aware & physics-aware AI systems.
Dataset Embodied Task Categories Data Domain Data Annotation Data Scale Spatial QAs
EmbSpatial-Bench [26] 6 Indoor (ScanNet, etc.) Template 2.2K 3.6K
Space3D-Bench [104] 6 Indoor (Replica) Manual 211 1K
Visual Spatial [61] 7 MSCOCO Template 10K 10K
SpatialRGPT-Bench [17] 12 Urban, Indoor, Sim Template 1.4K 1.4K
BLINK-Spatial [34] 14 MSCOCO Manual 286 286
What’s up [48] 6 Household, GQA, COCO Template 5K 5K
Spatial-MM [101] 4 Internet Template 2.3K 2.3K
RoboSpatial [102] 4 Indoor, tabletop Template 1M 3M
SpatialVLM [15] 2 WebLi Template 546 546
SpatialBot-Bench [11] 5 COCO,VG,RTX Manual 200 360
VSI-Bench [121] 8 Indoor Template 288 5K
OmniSpatial (Ours) 50 Internet Manual 1387 1.5K
Table 1: Comparison with other spatial reasoning benchmarks. A comparison between OmniSpatial and other existing spatial reasoning benchmarks reveals that, under the inclusion of Embodied Frames Compatibility, OmniSpatial avoids template-based annotations, features highly diverse data, and includes a significantly larger number of tasks.
The OmniSpatial benchmark includes images or video frames across diverse scenes, resolutions, lighting conditions, and weather patterns, collected from multiple countries across different continents. We evaluate state-of-the-art VLMs on our benchmark. Our findings indicate that, while current models perform well on conventional benchmarks, OmniSpatial presents a significantly greater challenge due to its comprehensive and complex task design.
Our key contributions are as follows:
  • We categorize visual-spatial reasoning into four key dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—broadening the scope of evaluation and guiding future research on spatial cognition in embodied & physics intelligence.
  • We develop the OmniSpatial dataset, which offers a diverse and challenging set of spatial tasks, serving as a comprehensive benchmark for assessing VLMs’ spatial reasoning capabilities.
  • We explore the enhancement of spatial reasoning in VLMs by incorporating auxiliary models in a chain-of-thought manner, demonstrating improved reasoning performance through this approach.
Figure 1: Benchmark Statistics of OmniSpatial: The distribution of tasks across 4 main categories.
2 Preliminaries: Visual–Spatial Reasoning
Spatial reasoning constitutes the cognitive bridge between visual perception and action planning in embodied agents. We define visual–spatial reasoning as the capacity of an artificial system to infer, predict, and manipulate spatial properties of the world from visual observations. Formally, let an RGB observation stream be $\mathbf{I}_{1:T}$ and a task-specific query be $q$. A model possesses visual–spatial reasoning if it can learn a mapping
$f:(\mathbf{I}_{1:T},\,q)\;\longrightarrow\;a,$
where $a$ belongs to a well-defined action or answer space whose correctness can be verified in the physical or simulated environment. This definition excludes non-visual priors so that improvements can be attributed to visual reasoning itself, yet it remains compatible with multi-modal extensions discussed in §1.
2.1 Taxonomy of Visual–Spatial Reasoning
Our classification stems from two complementary motivations. (i) Cognitive-psychology foundations. Decades of research on spatial cognition isolate partly independent faculties—e.g., spatial visualization, mental rotation, perspective taking, and spatial updating—that can be psychometrically evaluated in humans [13, 71]. These constructs offer a principled scaffold for analysing how agents perceive, reason about, and act within space. (ii) The need to move beyond basic spatial relations. Current spatial benchmarks are nearly saturated on tasks such as left–right discrimination, front–back identification, and object counting [15, 11, 17, 26, 104]. Real-world embodied tasks, however, require richer reasoning [50, 82, 16, 103, 53] about scene dynamics, multi-step spatial logic, physically grounded interaction, and viewpoint transformation.
Guided by these two strands, we partition spatial reasoning into four orthogonal yet complementary dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective taking. Each dimension aligns with a specific cognitive faculty and targets a distinct class of spatial challenges that is under-represented in prior work, enabling us to construct benchmarks that probe the full spectrum of complex spatial cognition while remaining anchored in established psychological theory.
Dynamic Reasoning involves predicting motion and temporal changes in spatial environments. This capability is crucial for tasks such as robotic control and autonomous navigation, where anticipating object trajectories ensures adaptive decision-making in dynamic contexts.
Complex Spatial Logic encompasses high-level reasoning about spatial relationships, transformations, and geometric structures. It is fundamental to problem-solving in domains such as engineering, design, and robotic manipulation, where understanding spatial configurations and predicting structural changes are essential.
Spatial Interaction focuses on engagement with spatial environments, where reasoning is guided by environmental constraints and task objectives. This includes applications such as path planning, obstacle avoidance, and real-time decision-making based on spatial feedback.
Perspective Taking refers to the ability to adopt different viewpoints, which is crucial for navigation, social cognition, and spatial awareness. It allows for understanding spatial relationships from multiple perspectives, facilitating decision-making in multi-agent environments and adaptive problem-solving.
2.2 Rationale for Classification
This taxonomy is designed to be both theoretically comprehensive and practically applicable. Dynamic reasoning emphasizes motion prediction, which is critical for navigation and interaction. Complex spatial logic provides a foundation for advanced cognitive tasks involving transformation and abstraction. Spatial interaction highlights real-time engagement with environments, while perspective taking captures cognitive flexibility in spatial reasoning. By structuring visual-spatial reasoning in this manner, we establish a framework that enhances the evaluation of spatial cognition models and their applications in artificial intelligence, robotics, and human cognition research.
Figure 2: Tasks Demonstration of OmniSpatial. Several representative subtasks are selected for demonstration in each of the four task categories. Note: The questions above are slightly simplified for clarity and conciseness.
Figure 3: Data Construction of OmniSpatial.The pipeline collects images from multiple sources and ensures their quality and relevance through manual selection. Precise annotations are then applied, ensuring each question has a clear, unique answer while maintaining a natural, conversational expression. This process supports the effective training of VLMs in spatial reasoning tasks.
3 OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark
3.1 Overview
We introduce OmniSpatial, a benchmark for the quantitative evaluation of VLMs in visual-spatial reasoning. Unlike large-scale benchmarks focused on a few tasks, OmniSpatial prioritizes task diversity and structured categorization over sheer dataset size. It consists of 1533 high-quality question-answer pairs, designed to challenge models beyond simple pattern matching or statistical learning, ensuring that spatial reasoning capabilities are tested rigorously, even with limited data.
Our dataset is curated from web searches, standardized tests, driving exam questions, and existing datasets including MME [57], HOI4D [67], etc.This diverse sourcing enhances the realism, complexity, and cross-domain generalization of the tasks. Web-sourced images cover natural environments, architecture, and daily life, adding visual complexity. Standardized tests from psychology and cognitive science provide scientifically rigorous spatial reasoning challenges. Driving exam questions introduce real-world dynamic interactions, such as road rule comprehension and motion prediction. Additionally, datasets like HOI4D contribute multi-resolution, varying illumination, and viewpoint diversity, as well as embodied intelligence tasks involving human-object interactions.
To ensure comprehensive evaluation, we categorize tasks into 4 major spatial reasoning types, further divided into 50 fine-grained task categories, as shown in Figs. 6 and 2. For example, perspective-taking tasks range from simple spatial judgments (“Is the opponent’s weapon to the left or right of the blue athlete?”) to complex motion prediction (“The blue athlete is blocking an incoming attack from the red athlete—what will happen next?”). Spatial interaction tasks include both static collision assessments (“Is the vehicle too close to the one in front?”) and dynamic environment reasoning (“A car door ahead is opening—a pedestrian exiting?”). This multi-level task structure ensures models are tested on diverse and progressively complex spatial reasoning challenges.
We emphasize precise image-task alignment in both dataset selection and annotation. Each question-answer pair is manually curated and reviewed in multiple rounds to ensure accuracy, consistency, and minimal ambiguity. Unlike automated annotation, manual refinement guarantees a high correlation between visual scenes and spatial reasoning tasks, strengthening OmniSpatial’s reliability as a benchmark for future research.
3.2 Benchmark Construction
This section describes how we collect data and annotations for OmniSpatial, as illustrated in Fig. 3.
3.2.1 Data Collection
As outlined in Section 2, our study defines 4 spatial categories with corresponding task types. To ensure a diverse and information-rich dataset, we design targeted search strategies for each task type, optimizing data collection for relevance, diversity, and complexity.
Web Images Web images constitute a major portion of our dataset, particularly for Perspective Taking, Dynamic Reasoning, and Spatial Interaction tasks. To ensure relevance and diversity, we formulate task-specific search terms reflecting distinct spatial scenarios, such as “indoor layout” and “furniture arrangement” for spatial relationship tasks. To minimize AI-generated image interference, we append filters like “-ai” and “-generated” to search queries. Using Google’s Custom Search JSON API, Web Robotic Process Automation (Web RPA), and manual searches, we retrieve thousands of images efficiently. A rigorous screening process is followed to eliminate irrelevant, low-resolution, or overly compressed images. We also removed images lacking spatial complexity, such as isolated static objects in Spatial Interaction tasks. The final dataset balances visual complexity and realism, enhancing its generalization capability across multiple tasks and scenarios while ensuring high-quality training data for VLMs. All the images are under MIT or CC-BY 4.0 licence.
Exam-Based Test Questions  Designed to assess complex spatial reasoning, these questions emphasize abstract spatial logic, including 3D transformations, rotations, and perspective shifts. We collect a wide range of publicly available spatial cognition tests via web scraping and manual curation. Noting variations in task focus, difficulty, and reasoning style, we categorized the questions to maintain a balanced representation. To prevent overrepresentation of any single reasoning type, we curate the dataset to ensure diverse question formats. Redundant, overly similar, or knowledge-intensive questions are removed, prioritizing those that purely test spatial reasoning. This refinement process significantly enhances the dataset’s quality and challenge level, providing a robust evaluation framework for VLMs.
Driving Test Questions  These questions evaluate spatial interaction in dynamic environments, focusing on distance estimation, right-of-way rules, and hazard detection. We source them from three primary channels: (i) Image-based multiple-choice questions from driving test websites in at least three countries, capturing diverse traffic rules and scenarios. (ii) Online question banks covering standardized spatial judgment tasks such as turning, lane changes, and parking decisions. (iii) Interactive driving videos from U.S. test resources, where we extract frames, annotate bounding boxes, and formulate contextual questions (“Which bounding box highlights a potential traffic risk?”). This multi-source approach ensures a realistic and challenging dataset, enhancing VLM adaptability in traffic-related spatial reasoning.
Existing Datasets  We integrate two key datasets: MME [57] and HOI4D [67]. MME provides RGB-D data, allowing depth-based spatial inference. We leverage its depth information for physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning. HOI4D contains extensive human-object interaction videos. We extract sequential frames to create motion prediction tasks, such as “Where will the hand holding the kettle move next?” By incorporating these datasets, we introduce real-world motion and interaction complexities, further strengthening VLMs’ dynamic reasoning capabilities.
Figure 4: PointGraph: Enhance Spatial Reasoning through Additional Scene Graphs.
Figure 5: SpatialCoT: Enhancing Spatial Imagination through Novel View Synthesis.
3.2.2 Question-Answer Annotation
We employ multiple-choice questions, including binary (true/false) and four-option formats, ensuring standardized evaluation and minimizing annotation errors. While adhering to structured formats, we prioritize natural and contextually rich language to prevent overly templated expressions.
Instead of rigid, mathematical phrasing (e.g., “Is [A] object on the right of [B] object?”), We use conversational styles more aligned with real-world reasoning (e.g., “If you’re entering the classroom, on which side are the students?”). This approach encourages models to rely on contextual and relative spatial references rather than predefined templates. To maintain clarity, we conduct multiple rounds of validation, resolving ambiguities in referents and ensuring unique answers. In particular, we ensure the quality of the data by means of cross-validation among different annotators.
We employ six human annotators for data labeling. All annotators receive comprehensive training to ensure a thorough understanding of the spatial reasoning tasks, enabling them to provide unbiased and discriminative assessments. All the annotators are from the authors of this work.
3.3 Improving Visual Spatial Reasoning Abilities
Given the diversity of spatial reasoning tasks in OmniSpatial, we further explore methods to enhance VLMs’ spatial reasoning capabilities.
3.3.1 PointGraph: Enhancing Spatial Reasoning via Point Relationships
We attempt to employ existing expert models, such as the Segment Anything Model [49], to partition images into distinct regions. This process extracts pixel clusters corresponding to multiple objects in the image, allowing models to estimate object centers and perform subsequent spatial reasoning tasks more effectively. As shown in Fig. 5, We attempt to enhance the spatial relationship comprehension of VLMs by explicitly incorporating additional spatial information cues into the models.
3.3.2 SpatialCoT: Enhancing Spatial Imagination via Novel View Synthesis
We observe that VLMs are largely constrained by textual reasoning paradigms, whereas humans inherently possess robust spatial imagination, commonly known as “mental imagery”. To some extent, this capability of spatial imagination corresponds to novel view synthesis of objects or scenes. Fortunately, recent advances in 3D generative models [100, 118, 87, 66] can perform such tasks effectively. In our work, we employ InstantMesh [118] to synthesize novel views from input images, providing additional visual cues to enhance spatial imagination.
Method Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
Blind Evaluation
Random Choice 24.98 - 24.86 26.30 25.88 23.43 27.27 21.44 24.77 22.55 24.84 25.78
GPT-3.5-turbo [98] 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
GPT-4-turbo [77] 34.06 - 42.97 37.40 41.18 28.95 40.00 22.27 26.32 31.37 33.99 35.42
Proprietary Models
GPT-4o-mini-2024-07-18 [45] 42.64 8 55.95 50.29 54.59 43.43 44.91 22.47 29.42 61.57 36.76 34.22
GPT-4o-2024-11-20 [45] 47.81 5 65.54 57.23 56.47 52.38 54.09 26.29 25.48 75.98 39.49 39.76
GPT-4.1-nano-2025-04-14 [78] 42.62 9 50.90 53.85 54.90 40.95 42.42 24.40 30.11 53.59 37.23 33.73
GPT-4.1-mini-2025-04-14 [78] 48.87 4 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
GPT-4.1-2025-04-14 [78] 51.78 2 66.22 64.74 60.00 65.33 60.18 31.75 30.06 70.98 40.64 39.04
Claude-3-5-sonnet-20241022 [3] 46.86 6 54.05 54.57 58.12 68.38 53.09 26.60 31.74 70.00 34.79 39.52
Claude-3-7-sonnet-20250219 [3] 47.53 5 57.57 55.95 56.71 63.81 59.09 29.48 28.39 72.16 36.06 36.63
Gemini-2.0-flash-lite-02-05 [2] 44.03 8 59.19 46.71 60.24 49.52 53.27 21.65 31.23 66.47 36.81 38.80
Gemini-2.0-flash-exp [2] 48.40 2 61.89 56.01 51.76 63.43 59.09 20.82 33.81 72.75 39.20 39.28
Gemini-2.5-flash-preview-05-20 [2] 52.12 1 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
Reasoning Models
o1-2024-12-17 [46] 50.36 6 71.62 60.98 57.65 63.81 60.00 39.18 27.10 71.57 38.03 36.14
o4-mini-2025-04-16 [79] 52.77 3 72.97 59.83 60.00 73.33 61.82 34.02 36.77 73.53 40.69 40.96
o3-2025-04-16 [79] 56.33 1 71.89 66.18 61.18 68.57 65.45 40.21 29.68 77.06 48.40 48.19
Claude-3-7-sonnet-20250219-thinking [3] 48.62 7 57.21 59.73 53.73 67.94 57.27 30.24 28.17 68.63 37.94 36.95
Gemini-2.5-flash-05-20-thinking [2] 53.16 3 70.27 64.74 61.18 72.38 58.18 35.05 36.13 74.12 40.96 32.53
Gemini-2.5-pro-preview-05-06 [2] 55.19 2 67.57 71.39 62.35 75.24 64.55 43.30 34.84 74.51 38.03 37.35
Open-source Models
LLavA-1.5-vicuna-7B [64] 34.97 15 54.46 31.23 35.29 36.19 33.94 29.01 24.18 55.60 34.66 36.14
LLaVA-onevision-qwen2-7B [54] 35.68 14 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
LLaVA-onevision-qwen2-72B [54] 45.66 6 62.16 50.29 54.12 60.95 56.36 22.68 25.81 76.47 37.23 33.73
Gemma-3-4B [47] 39.79 11 41.89 49.71 56.47 27.62 36.36 23.71 24.52 59.80 36.17 38.55
Gemma-3-12B [47] 43.71 8 54.05 54.91 54.12 47.62 45.45 16.49 30.32 63.73 36.70 33.73
Gemma-3-27B [47] 44.75 7 56.76 55.78 57.65 50.48 52.73 27.84 29.03 64.71 33.51 32.53
InternVL3-2B [130] 37.98 13 50.00 40.58 43.29 40.00 40.55 21.86 28.52 55.49 35.11 33.01
InternVL3-8B [130] 41.60 9 52.43 40.87 48.94 51.05 44.77 24.95 28.63 64.20 38.62 40.96
InternVL3-14B [130] 45.94 5 54.32 60.17 50.35 51.81 51.45 28.04 28.26 68.04 35.37 34.46
InternVL3-38B [130] 48.48 2 63.42 63.58 54.59 58.29 50.55 29.90 28.52 72.16 36.76 33.49
InternVL3-78B [130] 49.33 1 63.78 63.12 56.24 59.24 51.45 27.63 30.19 74.51 38.46 35.90
Qwen-VL2.5-3B [115] 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
Qwen-VL2.5-7B [115] 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
Qwen-VL2.5-32B [115] 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
Qwen-VL2.5-72B [115] 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
Specialized Spatial Reasoning Models
SpaceMantis-13B [15] 36.36 6 47.03 36.59 40.94 34.86 33.09 22.27 24.39 49.22 38.25 39.28
SpaceQwen2.5-VL-3B [15] 40.25 3 58.11 39.88 41.18 40.95 40.91 29.90 25.81 63.73 38.83 39.76
SpaceThinker-Qwen2.5VL-3B [15] 40.42 2 47.84 53.06 43.29 35.43 38.73 24.33 28.00 58.04 35.11 31.08
SpatialBot-3B [11] 35.68 6 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
RoboPoint-vicuna-v1.5-7B-lora [123] 35.85 6 57.03 28.61 34.82 37.33 40.55 29.90 22.71 50.20 38.72 40.96
RoboPoint-vicuna-v1.5-13B [123] 34.60 5 55.68 28.15 42.82 32.19 32.55 24.12 27.74 49.02 37.66 33.49
SoFar-Qwen2.5VL-3B [89] 45.14 1 56.49 51.16 54.12 53.14 52.73 31.75 22.88 71.60 36.56 41.69
Human Evaluation
Human 92.63 - 96.53 97.30 92.94 97.14 94.55 91.30 87.63 99.02 95.74 93.98
Table 2: Evaluation on OmniSpatial. All models were tested 5 times and averaged to reduce randomness. Dark green indicates the best result within the group, and light green indicates the second-best result within the group. Gemini-2.5-Pro [97], InternVL3-78B [130], and SoFar [89] achieve the best performance within their respective groups. Among the various spatial intelligence categories, Complex Logic is the most challenging.
4 Experiments
4.1 Evaluation Setup
To assess the spatial reasoning capabilities of existing VLMs, we conduct comprehensive experiments on OmniSpatial. This evaluation compares both proprietary and open-source models across various spatial tasks, using standardized benchmarks to ensure fairness and reproducibility.
Benchmark Models  We evaluate a range of state-of-the-art vision-language models, including:
  • Proprietary Models: These include GPT-4o series [45], GPT-4.1 series, Claude series [3], and Gemini series [2], which are state-of-the-art multimodal models accessed via an API under zero-shot settings with standardized prompts. All the system prompts can be found at  Appendix D.
  • Reasoning Models: We refer to current models employing internal long CoT as reasoning models. These models typically leverage reinforcement learning to enhance their reasoning capabilities, including models such as the o1, o3, o4-mini [46, 79], and Gemini-2.5 series [2, 97]. Due to the explicit nature of CoT, we utilize an additional text-based evaluation model [98] to determine the correctness of responses, similar to LLM-as-a-Judge [129].
  • Open Source Models: Our open-source set comprises Qwen-VL series [6], InternVL series [130], Gemma series [47] and LLaVA-Onevision series [63]. These models are currently the most advanced open-source VLMs. All models are deployed locally and use standardized prompts.
  • Specialized Spatial Models: These models, including SpatialVLM [15], RoboPoint [123], SpatialBot [11], and SoFar [89], are designed specifically for spatial reasoning tasks. SpatialVLM integrates metric spatial information through 3D data synthesis and scene segmentation. RoboPoint and SpatialBot emphasize the role of point affordance in spatial reasoning.
4.2 Evaluation Metrics
We evaluate the spatial reasoning capabilities of models by measuring their accuracy on multiple-choice questions. For standard open-source and closed-source models, we experiment with four approaches: enforcing direct output result, regular expression parsing, JSON format parsing and LLM-as-a-Judge [129]. For Reinforcement Learning-based reasoning models, due to the inherent challenges in enforcing structured output, we employ GPT-4.1-mini to assess the correctness of the model’s responses by comparing them with the ground truth. More detailed ablation studies are discussed in the Section C.1.
Method Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
GPT-4.1-mini - - - - - - - - - - - -
(w/o CoT) 48.86 - 64.05 58.55 57.65 59.43 56.91 28.87 34.06 68.82 37.18 41.20
(w/ Zero-shot CoT) 49.81 +0.95 62.97 58.96 59.06 62.48 58.55 27.63 32.52 69.41 40.11 40.96
(w/ Manual CoT) 49.76 +0.90 65.68 58.90 58.59 64.38 56.91 28.45 32.13 69.61 39.31 41.20
(w/ PointGraph) 50.49 +1.63 67.57 62.14 57.65 64.76 58.18 28.87 30.32 70.59 38.83 42.17
Gemini-2.5-Flash - - - - - - - - - - - -
(w/o CoT) 51.47 - 66.22 65.90 63.53 71.43 66.36 32.99 34.84 70.59 31.91 38.55
(w/ Zero-shot CoT) 51.53 +0.06 63.51 61.27 58.82 67.62 65.45 42.27 34.84 79.41 35.90 32.53
(w/ Manual CoT) 52.12 +0.65 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
(w/ PointGraph) 53.23 +1.76 62.16 69.94 64.71 67.62 59.09 29.90 38.06 74.51 37.77 37.35
Qwen-VL2.5-3B - - - - - - - - - - - -
(w/o CoT) 41.45 - 58.65 43.06 39.53 50.67 48.73 32.78 22.58 61.96 37.66 37.35
(w/ Zero-shot CoT) 40.64 -0.81 59.73 43.87 46.12 48.38 43.27 25.36 22.84 59.61 36.54 37.59
(w/ Manual CoT) 40.07 -1.38 55.68 46.65 47.29 40.57 46.00 28.04 24.39 60.39 33.35 31.57
(w/ PointGraph) 44.36 +2.91 55.68 55.20 48.94 52.19 52.36 29.90 25.55 66.08 35.11 31.08
Table 3: Camparison of textual Chain-of-Thought and PointGraph on OmniSpatial.
4.3 Main Results
Overall Model Performance  As illustrated in Table 2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 [79] and Gemini-2.5-pro [2], achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and require a lot of inference time and tokens. (ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B [130] and Qwen-VL2.5-72B [115] achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp. (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
Category-wise Analysis  Breaking down the results, we observe notable performance differences across spatial reasoning categories: (i) Leveraging their extensive world knowledge and local understanding capabilities, proprietary models have demonstrated strong performance in Dynamic Reasoning and Spatial Interaction, indicating that reasoning models possess high proficiency in temporal understanding, spatial relationship analysis, and map-based comprehension. (ii) For Pattern Geometric and Recognition Reasoning, which involve spatial imagination in planar geometry, even reasoning models designed for extended thinking can only achieve an accuracy of around 30% to 40%, slightly surpassing the random baseline. (iii) Current models exhibit limited Perspective Taking abilities, predominantly analyzing scenarios from an ego-centric viewpoint while struggling to imagine perspectives from others’ viewpoints.
Method Avg. Improve
Ego
Centric
Allo
Centric
Hypothetical
GPT-4.1-mini
(w/ Zero-shot CoT) 45.56 - 69.41 40.11 40.96
(w/ Spatial CoT) 47.58 +2.02 69.43 42.37 44.34
Qwen-VL2.5-3B
(w/ Zero-shot CoT) 40.89 - 59.61 36.54 37.59
(w/ Spatial CoT) 42.90 +2.01 60.80 39.25 37.44
Table 4: Performance of Spatial CoT on OmniSpatial Perspective-Taking track.
Impact of PointGraph & Spatial CoT  To test whether structured segmentation improves performance, we apply PointGraph as a preprocessing step for GPT-4.1, Gemini-2.5-flash and Qwen-VL2.5-7B. Results in Table 3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement. Fig. 5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track. Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.
5 Related Works
5.1 Benchmarking Spatial Reasoning
The development of standardized benchmarks [55, 122, 65, 124, 44, 51] has been critical in assessing the performance of spatial reasoning models. Various studies have introduced innovative benchmarking methodologies [104, 15, 17, 11, 102, 89, 96] to advance spatial reasoning evaluation. For instance, Spatial VQA [26] was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning. SpatialRGPT-Bench [17] further introduced depth-aware reasoning, utilizing large language models to analyze spatial relationships, such as understanding relative directions and distances between specified regions in 3D scenes [104]. SpatialBench [11] categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks. RoboSpatial [102] proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives. VSI-Bench [121] combined video data [14, 65, 33, 56, 32] with cognitive maps [72, 4] to simulate human-like spatial cognition and optimize reasoning in dynamic environments. Recently, SoFar [89] introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking. Inspired by prior spatial reasoning research [119, 114, 58, 76], we identified several limitations in existing benchmarks, such as reliance on generated images [104, 34, 48, 61, 92, 101], LLM-generated templates [59, 26, 91, 93], and domain-specific focus [117, 19, 21, 12, 28]. These issues hinder their comprehensiveness and real-world applicability. To address these gaps, our study proposes a comprehensive and integrative spatial reasoning benchmark.
6 Conclusion
In this paper, we introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning. OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 1.5 K manually-curated question–answer pairs. Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking. By exposing these systematic gaps, OmniSpatial provides a rigorous yardstick and a rich error analysis playground for future work on physics- and viewpoint-aware multimodal models.
Appendix A Detailed Task Design
OmniSpatial aims to comprehensively evaluate the spatial reasoning capabilities of Vision-Language Models, covering four major categories: Dynamic Reasoning, Complex Logic, Spatial Interaction, and Perspective Taking. Each category not only focuses on different types of spatial reasoning tasks but also includes challenges based on real-world application scenarios. This approach helps researchers better understand and enhance models’ multi-domain spatial reasoning abilities. The following presents the underlying considerations and practical value behind each task design.
A.1 Dynamic Reasoning
The Dynamic Reasoning category focuses on the model’s understanding of object movement and its changes, assessing the ability to make accurate judgments in uncertain or rapidly changing environments. Spatial dynamics are critical not only in robot control but also have broad applications in fields like autonomous driving and intelligent surveillance.
A.1.1 Manipulation
Operational Position Selection  This task evaluates how models determine the optimal interaction point with objects in complex environments. Selecting the best grasping point can prevent tilting or damage to objects and improve the efficiency and precision of operations. This task is crucial in robotic grasping, especially when environmental conditions are unstable.
Movement Direction Determination  This task assesses the model’s ability to predict the movement direction of an object or itself, providing decision support for automated systems and helping to optimize robot motion strategies.
Intent Recognition  Intent recognition involves inferring the purpose or goal behind a movement. This task is particularly important for contextual analysis, such as determining whether a person is reaching for a door handle to open or close a door. This capability enhances the reasoning of human-robot interactions and optimizes the interaction experience of intelligent assistants and robots.
A.1.2 Motion Analysis
Uniform Motion  The model’s ability to reason about uniform motion reflects its fundamental understanding of time and spatial relationships, such as estimating the speed or time required for a target to move. This is applicable in areas like object tracking and path prediction, such as estimating vehicle travel time or train arrival time.
Variable Motion  Variable motion analysis focuses on understanding acceleration and deceleration processes. By predicting the position changes of an object during variable motion, the model can better simulate dynamic phenomena in the physical world, such as calculating braking distance for vehicles. This is critical in autonomous driving and robot control.
Spatial Compatibility  Tasks that assess whether an object fits within a specific space directly relate to the precision of robots and automated devices in real-world operations. For instance, determining whether luggage can fit into an overhead compartment is applicable in logistics and automated warehouses, helping systems make effective object adaptation decisions.
A.2 Complex Logic
The Complex Logic category focuses on higher-order spatial reasoning, including tasks such as geometric transformations and pattern recognition, challenging the model’s ability to abstractly understand and reason in multi-dimensional spatial environments.
A.2.1 Pattern Recognition
Style  Style recognition tasks assess the model’s ability to infer visual rules in structured patterns. Typical operations include completing missing parts, combining or subtracting shapes, comparing similarities and differences, and performing visual logic like black-white inversion. This skill is crucial for tasks like intelligence tests and diagrammatic reasoning.
Quantity  Quantity-based tasks evaluate the model’s ability to perform visual numerosity reasoning, focusing on implicit quantitative patterns such as the number of points, lines, regions, elements, or strokes. These tasks require abstract counting under diverse spatial configurations, often without explicit numerical labels or symbols.
Attributes  Attribute-based tasks evaluate the model’s ability to reason about non-numeric visual properties such as symmetry (axial or radial), curvature, and openness. These tasks require the recognition of structural features that do not rely on quantity but rather on geometric or perceptual traits.
Location  Location tasks evaluate the model’s ability to reason about spatial changes such as translation, rotation, and reflection. The focus is on how objects shift in space while preserving their structure. This skill is essential for grid-based reasoning, motion prediction, and geometric pattern understanding.
A.2.2 Geometric Reasoning
Polyhedron Unfolding  This task examines whether the model can infer the 2D net of a 3D object, reflecting its ability to mentally construct spatial layouts. It is particularly relevant to applications such as packaging design, industrial manufacturing, and aerospace engineering.
Sections and Projections  By evaluating how models interpret 2D cross-sections or projections from different viewpoints, this task challenges their capacity to connect visual appearance with internal structure—key to fields like medical imaging, architecture, and mechanical design.
Mental Rotation  Mental rotation requires the model to simulate the rotation of objects in mind and track changes across views. It underpins spatial reasoning in tasks ranging from CAD modeling to object manipulation in virtual and augmented reality.
Assembly  This task involves reasoning about how separate parts fit together into a coherent whole, testing the model’s understanding of geometric constraints. It has broad relevance in robotics, structural analysis, and physical assembly planning.
Analytical Geometry  These tasks are inspired by classic geometry problems, requiring models to reason about spatial relations using angles, distances, and symmetry. They bridge mathematical logic with visual structure, supporting applications in structured reasoning and spatial abstraction.
A.3 Spatial Interaction
The Spatial Interaction category evaluates a model’s ability to reason about interactions with objects and environments. It includes tasks such as Traffic Analysis, Localization, and Geospatial Strategy, reflecting the model’s understanding and application of spatial knowledge in real-world scenarios.
A.3.1 Traffic Analysis
Anomaly Detection  This task focuses on identifying potential dangers or traffic violations in complex scenes, such as unsafe following distances or unusual vehicle behavior. It plays a key role in ensuring safety in autonomous driving systems.
Sign Recognition  This task evaluates the model’s ability to detect and interpret traffic signs, including speed limits, no-entry zones, and yield signs. Accurate recognition is critical for safe and rule-compliant decision-making.
Action Recognition  This task involves identifying or predicting the actions of traffic participants, such as driver gestures, police signals, or pedestrian intentions. It is important for understanding dynamic human behavior in traffic environments.
Risk Detection  This task aims to detect immediate hazards in the environment, such as an opening car door or a pedestrian crossing the road. Timely detection supports effective avoidance and control strategies.
Behavior Guidance  This task provides context-aware behavioral suggestions, such as advising to turn off high beams or reminding that parking is prohibited. It enhances overall driving safety and compliance.
Contextual Analysis  This task assesses the model’s ability to interpret spatial relations and behaviors based on environmental cues. For example, estimating wind conditions to anticipate overtaking behavior or understanding road status to infer potential risks.
A.3.2 Localization
UI Interaction  This task requires the model to determine which icon should be selected within a user interface based on contextual understanding, and to accurately localize its position. It reflects the model’s ability to integrate semantic interpretation with spatial reasoning, supporting applications in intelligent assistants and automated interface control.
Object Detection  This task involves identifying specific target objects within an image. It is often paired with spatial localization to jointly assess what the object is and where it is located.
Spatial Localization  This task focuses on determining the precise position of objects within a scene. It is commonly evaluated alongside object detection to answer questions like “What is at this location?” or “Where is this object?”
Pose Estimation  This task estimates the orientation and spatial configuration of objects, such as detecting whether a cup is upright or tipped over. It is frequently integrated with spatial localization to enable more nuanced scene understanding.
A.3.3 Geospatial Strategy
Spatial Recognition  Assesses the model’s capacity to identify spatial structures such as rooms, corridors, or zones within a scene. This is essential for semantic navigation and indoor mapping.
Location Recognition  This task evaluates the model’s ability to identify specific locations on a map or scene, such as recognizing the position marked as “You are here” or locating a designated landmark. It reflects the model’s capacity to associate spatial markers with real-world positions.
Region Recognition  Focuses on distinguishing and classifying regions in a broader spatial context, such as residential vs. industrial zones on a map.
Route Interpretation  Tests the model’s ability to follow or explain a route depicted in a map or scene. It requires understanding directional arrows, route labels, and spatial transitions.
Route Design  Involves selecting or generating an optimal path to reach a given goal, considering spatial constraints and possible alternatives.
Route Selection  Compares multiple candidate routes and chooses the most suitable one based on efficiency, safety, or contextual requirements.
Navigation  Evaluates the model’s ability to understand smartphone or in-vehicle navigation interfaces, including interpreting turn-by-turn directions, identifying route segments, and understanding map overlays. This is crucial for building intelligent voice assistants and real-time guidance systems.
Map/Scene Conversion  Tests the ability to mentally convert between map views and real-world scenes, which is critical in correlating schematic representations with physical surroundings.
Legend Recognition  It requires identifying and interpreting map symbols (e.g., stairs, elevators, emergency exits) and a foundational skill in navigation and spatial reasoning.
Terrain Identification  Focuses on distinguishing types of terrain (e.g., flat, uphill, water-crossing), which is essential for planning safe and feasible paths in outdoor navigation or robotics.
A.4 Perspective Taking
This category evaluates the model’s ability to understand spatial relationships from different viewpoints. Since changes in perspective directly affect what is observed, the ability to reason across varying angles is essential for robotic perception and interaction.
A.4.1 Egocentric
Count  Counting the number of visible objects from the current perspective is crucial for dynamic interaction. For example, in robotic grasping tasks, knowing how many targets are visible helps determine the appropriate operation strategy.
Size  Judging the size of an object from the observer’s viewpoint aids robots or virtual systems in depth perception, helping assess whether an object can be grasped or properly placed.
Direction  This refers to the direction directly seen by the observer. It is especially important in autonomous driving scenarios, where understanding object movement helps predict traffic conditions and enables timely responses.
Order  Analyzing the arrangement of multiple objects in an image is essential for robotic operations, helping prioritize which objects to interact with first.
Distance  The distance between objects as perceived from the observer’s viewpoint is a key capability in navigation systems, enabling path planning and obstacle avoidance.
A.4.2 Allocentric
Count  Understanding how the perceived quantity of objects changes under different viewpoints is key. Due to variations in position and orientation, occlusion often occurs—for example, a driver may have blind spots, while a road surveillance camera can see objects hidden in those areas. This task evaluates the model’s ability to judge the difference in object counts from various observation points.
Size  This task involves evaluating an object’s size from a specified observer’s viewpoint. Due to the general principle that closer objects appear larger and distant ones appear smaller, objects of the same size may look different to different observers. The model is expected to either infer the true size based on reference objects or estimate how perceived size changes with viewing position.
Direction  This task emphasizes judging directions from abstract positions, such as another agent’s viewpoint or a map. The answers often differ from what is directly observed, requiring one to adopt the target’s perspective—engaging in perspective-taking. It is crucial not only for large-scale path planning but also for understanding an object’s intrinsic orientation.
Order  This task requires observing the arrangement of objects from a specified viewpoint—for example, the seating order of students in the front row as seen by a teacher on the podium, which is exactly reversed from what a camera at the back of the classroom would capture. Only by understanding “what the target sees” can one make accurate predictions or judgments about the scene.
Distance  Differences in the observer’s position and orientation lead to variations in perceived size and distance. For example, in the same scene, a photographer taking pictures of the same object from different camera angles will capture different impressions of its size and distance. Changing the camera position essentially means changing the perspective. This ability is vital for coordination and environmental assessment in human-robot interaction or multi-robot collaborative tasks.
A.4.3 Hypothetical Perspective Taking
This category focuses on imagining the scene from a specified but non-existent viewpoint, requiring the model to mentally adopt a fictional position—an advanced form of perspective-taking in spatial reasoning.
Count  Predicting how many target objects would be visible from a hypothetical viewpoint—for example, a person standing at the opposite corner of the street may see a different distribution of objects. The model must reason about occlusion, orientation, and visibility from the imagined perspective.
Size  Inferring the apparent size of objects from a hypothetical location and direction. For instance, the same object may appear larger when viewed up close or smaller when viewed from above. The model needs to simulate how visual scale changes with altered viewpoints.
Direction  Reasoning about how an object’s direction appears from a location where no observer is present. For example, a pedestrian walking toward a doorway would appear “head-on” from the entrance, but present a different direction from a side view.
Order  Simulating the arrangement of objects from another location helps assess how spatial sequences change across viewpoints. For example, the seating order seen from the podium may be the reverse of what’s seen from the back of the room.
Distance  Estimating relative distances between objects from a hypothetical position requires mentally adopting a new viewpoint. This supports effective planning and coordination in tasks such as multi-robot navigation or collaborative manipulation.
Appendix B Additional Related Works
B.1 Spatial Vision-Language Models
Spatial vision-language models integrate computer vision [1, 37, 80, 90] and natural language processing [6, 115, 9, 2, 107, 108, 128] to enhance machine understanding of spatial relationships. Recent research [88, 15, 17, 11, 102, 123, 89] has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments. One of the pioneering works in this domain, SpatialVLM [15], has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset. These models [45, 45, 97, 115] effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
Further advancements [62, 60, 94, 105, 120, 99, 116, 103] include SpatialRGPT [17], which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities. Similarly, SpatialBot [11] explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments. Recently, SoFar [89] proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities [18, 25, 73, 131, 113, 43, 74, 27, 42, 89]. However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks. To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
B.2 Spatial Reasoning in Psychology
In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings [35]. To systematically characterize this ability, Buckley et al. [10]proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation. Malanchini et al. [69] further identified a strong correlation between spatial orientation and object manipulation skills. Additionally, Newcombe & Shipley [75]introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic. Furthermore, many studies also have contributed to the development of frameworks [112, 52, 69, 39, 70] and evaluation methods [29, 111, 110, 40] for spatial reasoning. Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning. These psychological frameworks provide complementary perspectives for understanding and assessing spatial reasoning, offering theoretical foundations for embodied intelligent systems. Inspired by these classifications, our study proposes a novel taxonomy for visual-spatial reasoning, aiming to advance spatial reasoning research in vision-language models.
Appendix C Additional Ablation Studies
C.1 Evaluation and Format Strategy
The choice of evaluation format and answer-extraction strategy can materially influence how fully a large language model’s capabilities are expressed. Because our benchmark introduces a novel visual–spatial reasoning task, Table 5 reports the performance impact of different prompting and evaluation schemes. We consider three prompting paradigms—Direct QA, Zero-shot CoT, and Manual CoT—and three evaluation methods: (i) regex-based field extraction (RE), in which the VLM is instructed to place its answer in a fixed slot at the end of the response; (ii) JSON extraction, where the answer must appear in a dedicated answer field; and (iii) LLM-based free-form evaluation, for which we employ GPT-4.1-mini as an adjudicator.
We observe that reasoning-centric models that strongly obey formatting directives (e.g., Gemini-2.5-flash) cannot be reliably scored with regex extraction, because their compulsory chain-of-thought interferes with the required output template. For standard models, all three evaluation strategies yield comparable scores, and both CoT variants consistently outperform the Direct QA baseline by a small margin. Accordingly, in the main results table we evaluate non-reasoning models with the Manual CoT + RE setting, whereas reasoning-oriented models are assessed with Manual CoT + LLM evaluation.
Promopt Type Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
Manipulate
Motion
Analysis
Traffic
Analysis
Locate
Geospatial
Strategy
Pattern
Recognition
Geometric
Reasoning
Ego
Centric
Allo
Centric
Hypothetical
GPT-4.1-mini - - - - - - - - - - - -
None Direct 39.22 57.57 36.18 58.12 44.95 53.45 12.58 18.58 62.94 37.61 37.83
None RE 48.86 64.05 58.55 57.65 59.43 56.91 28.87 34.06 68.82 37.18 41.20
Zero-shot CoT RE 49.81 62.97 58.96 59.06 62.48 58.55 27.63 32.52 69.41 40.11 40.96
Manual CoT RE 49.76 65.68 58.90 58.59 64.38 56.91 28.45 32.13 69.61 39.31 41.20
None JSON 48.23 61.08 55.55 57.41 58.86 54.00 28.89 29.16 70.20 40.96 40.48
Zero-shot CoT JSON 48.70 58.67 57.17 57.88 57.71 55.09 28.89 29.29 69.02 40.16 42.89
Manual CoT JSON 48.87 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
None LLM 48.02 60.54 56.82 58.59 58.10 57.45 28.04 32.90 68.24 37.55 38.31
Zero-shot CoT LLM 48.36 64.05 56.71 58.59 58.86 57.27 27.84 33.29 67.06 38.30 38.80
Manual CoT LLM 49.85 62.97 59.48 58.12 61.33 58.36 28.04 31.61 69.02 41.12 39.28
Gemini-2.5-flash - - - - - - - - - - - -
None LLM 51.47 66.22 65.90 63.53 71.43 66.36 32.99 34.84 70.59 31.91 38.55
Zero-shot CoT LLM 51.53 63.51 61.27 58.82 67.62 65.45 42.27 34.84 79.41 35.90 32.53
Manual CoT LLM 52.12 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
Qwen-VL2.5-3B - - - - - - - - - - - -
None Direct 44.04 60.27 52.20 47.53 50.48 52.73 22.68 30.32 65.49 36.49 30.84
None RE 41.45 58.65 43.06 39.53 50.67 48.73 32.78 22.58 61.96 37.66 37.35
Zero-shot CoT RE 40.64 59.73 43.87 46.12 48.38 43.27 25.36 22.84 59.61 36.54 37.59
Manual CoT RE 40.07 55.68 46.65 47.29 40.57 46.00 28.04 24.39 60.39 33.35 31.57
None JSON 38.08 62.97 32.49 46.59 48.95 47.64 24.54 23.61 62.16 35.85 27.47
Zero-shot CoT JSON 39.20 61.35 40.40 48.94 45.52 47.27 22.27 24.00 65.29 33.51 27.71
Manual CoT JSON 38.37 58.11 34.28 42.35 46.29 48.55 27.63 24.65 62.75 36.54 26.75
None LLM 42.10 66.22 43.99 44.00 51.24 50.73 27.84 25.42 70.00 35.90 29.40
Zero-shot CoT LLM 35.89 54.59 36.24 49.41 40.19 40.18 24.95 20.13 61.37 29.20 33.98
Manual CoT LLM 36.26 53.78 34.10 50.35 42.10 44.00 23.51 26.06 60.39 31.38 23.86
Table 5: Comparative analysis of various prompting and evaluation strategies.
Appendix D System Prompts
We present all the system prompts used in our experiments in Figs. 7 and 8 to facilitate reproducibility. We observe that some models are sensitive to the choice of system prompt, which may stem from distributional biases in their training data. In contrast, inference-oriented models generally exhibit stronger generalization capabilities and are less reliant on carefully crafted prompts.
Appendix E Additional Visualization
We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs. 9, 10, 11 and 12, respectively. Our benchmark comprises a rich collection of data samples spanning diverse scenarios, resolutions, lighting conditions, and geographical regions. It includes both absolute numerical analyses and relative spatial relationships, aiming to comprehensively evaluate the spatial reasoning capabilities of vision-language models.
Appendix F Limitation & Future Works
Although OmniSpatial includes some image clips with dynamic information from HOI4D [67], the complexity of the operational tasks still lags behind that of long videos [121, 14]. Moreover, while PointGraph & Spatial CoT enhances VLM’s spatial understanding through point cues, the improvement is not fundamental in nature. Spatial reasoning tasks are more like mathematics and coding tasks, require longer and more complex reasoning [22, 46].
3D information is crucial for spatial reasoning, and future work involves introducing 3D representation [23, 86, 88, 87] and perception [84, 85, 127, 30], as well as 3D VLMs [81, 24, 36, 88, 89], reasoning model [125, 22, 79] and knowledge distillation [126, 41, 23, 86]. The ultimate goal of spatial reasoning is to empower robots, and future work also involves robot execution tasks [42, 31, 89, 43, 38].
Appendix G Broader Impacts
OmniSpatial promises several positive societal benefits. By pushing models to reason about motion, collision risk and traffic scenes, it can hasten the arrival of safer autonomous vehicles and service robots that foresee hazards and navigate crowded spaces responsibly. Its geometric-reasoning tasks—from polyhedron unfolding to assembly—offer data that can streamline product design, packaging and manufacturing, lowering material use and energy waste. The benchmark’s localization, UI-interaction and perspective-taking challenges cultivate spatially aware assistants that improve AR/VR experiences, access tools for visually impaired users and more natural human-computer interfaces.
Figure 7: System prompts used in OmniSpatial evaluation.
Figure 8: System prompts used in OmniSpatial evaluation.
Figure 9: Visualization example of OmniSpatial data samples.
Figure 10: Visualization example of OmniSpatial data samples.
Figure 11: Visualization example of OmniSpatial data samples.
Figure 12: Visualization example of OmniSpatial data samples.

### v2

@@
 Abstract
-Spatial reasoning is a key aspect of cognitive psychology and remains a major bottleneck for current vision-language models (VLMs).
-While extensive research has aimed to evaluate or improve VLMs’ understanding of basic spatial relations, such as distinguishing left from right, near from far, and object counting, these tasks represent only the most fundamental level of spatial reasoning.
+Spatial reasoning is a key aspect of cognitive psychology and remains a bottleneck for current vision-language models (VLMs).
+While extensive research has aimed to evaluate or improve VLMs’ understanding of basic spatial relations, such as distinguishing left from right, near from far, and object counting, these tasks cover only the most elementary layer of spatial reasoning and are largely approaching saturation in the latest reasoning models.
 In this work, we introduce OmniSpatial, a comprehensive and challenging benchmark for spatial reasoning, grounded in cognitive psychology.
 OmniSpatial covers four major categories: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, with 50 fine-grained subcategories.
-Through Internet data crawling and careful manual annotation, we construct over 1.5K question-answer pairs.
-Extensive experiments show that both open- and closed-source VLMs, as well as existing reasoning and spatial understanding models, exhibit significant limitations in comprehensive spatial understanding.
-We further analyze failure cases and propose potential directions for future research.
+Through careful manual annotation, we construct over 8.4K question-answer pairs.
+Extensive experiments show that both open- and closed-source VLMs exhibit significant limitations in comprehensive spatial reasoning.
+We also explore two strategies—PointGraph (explicit scene graph cues) and SpatialCoT (novel-view chain-of-thought)—to bolster spatial reasoning.
 Overview of OmniSpatial Benchmark.
 1 Introduction
-Spatial reasoning plays a crucial role in bridging visual observation to robotic action [43, 88, 89], autonomous driving, and AR/VR.
+Spatial reasoning plays a crucial role in bridging visual observation to robotic action (Huang et al., 2024b, Qi et al., 2024; 2025), autonomous driving, and AR/VR.
 For models to execute tasks effectively, they must understand spatial relationships to determine appropriate actions.
-To enhance the spatial understanding of Vision-Language Models (VLMs), prior works [15, 17, 11, 68, 102, 123, 121, 89] have integrated spatial information into datasets, enabling basic forms of spatial reasoning.
-Various benchmarks [26, 104, 101, 48, 34, 102, 121, 89] have been introduced to systematically evaluate such capabilities, focusing on tasks like recognizing left and right, estimating depth, and constructing cognitive maps [121, 106, 72].
-Additionally, spatial reasoning has been applied to manipulation tasks [25, 8, 131, 20, 123], allowing systems to position objects according to specified spatial rules [89].
-Despite these advances, existing approaches largely focus on basic spatial understanding, which typically includes identifying simple spatial relations such as relative positions (left, right, front, back), proximity (near, far), and object-level counting.
-Recent models, including state-of-the-art reasoning systems like ChatGPT o3 [79] and Gemini-2.5-Pro [97], have achieved over 90% accuracy on benchmarks such as SpatialBot-Bench [11] and EmbSpatial [26], suggesting that these basic tasks are approaching saturation, as shown in OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models.
-However, complex spatial reasoning remains a significant challenge [35, 5, 83, 50, 82, 16].
-Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships [7, 109, 95].
+To enhance the spatial understanding of Vision-Language Models, prior works (Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Ma et al., 2024, Song et al., 2024, Yuan et al., 2024, Yang et al., 2024, Qi et al., 2025) have integrated spatial information into datasets, enabling basic forms of spatial reasoning.
+Various benchmarks (Du et al., 2024, Szymanska et al., 2024, Shiri et al., 2024, Fu et al., 2024b, Song et al., 2024, Yang et al., 2024) have been introduced to systematically evaluate such capabilities, focusing on tasks like recognizing left and right, estimating depth, and constructing cognitive maps (Yang et al., 2024, Tolman, 1948, Momennejad et al., 2023).
+Additionally, spatial reasoning has been applied to manipulation tasks (Driess et al., 2023, Qi et al., 2025, Yuan et al., 2024), allowing systems to position objects according to specified spatial rules.
+However, existing benchmarks still target basic spatial understanding, such as position relationship (left, right, front, back), proximity (near, far), and object counting.
+We use the latest reasoning models and agents to evaluate these benchmarks, such as o3 (OpenAI, 2025b) and Gemini-2.5-Pro (Reid et al., 2024).
+The results are shown in the lower left of OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models.
+These models have achieved over 90% accuracy on previous benchmarks such as SpatialBot-Bench (Cai et al., 2024) and EmbSpatial (Du et al., 2024), suggesting that these basic tasks are approaching saturation.
+We believe complex spatial reasoning remains a significant challenge (Gardner, 2011, Baddeley, 1998, Previc, 1998, Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025).
+Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships (Bar-Anan et al., 2006, Trope & Liberman, 2010, Ramalho et al., 2018).
 For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route.
 Similarly, tasks like inserting a knife into a rack or flattening a box demand reasoning about object rotation, deformation, and spatial compatibility—far beyond static object placement.
-From the perspective of cognitive psychology, complex spatial reasoning extends beyond basic relations and includes dynamic world knowledge reasoning, interactive spatial behavior with environments or agents, logical analysis of 3D spatial structures, and perspective-taking abilities.
-Therefore, to address these limitations and systematically assess higher-order spatial reasoning, we propose OmniSpatial, a comprehensive benchmark capturing the breadth and depth of spatial cognition.
-OmniSpatial categorizes spatial reasoning into four key dimensions: dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking, providing a foundation for advancing the next generation of spatially-aware & physics-aware AI systems.
+From the perspective of cognitive psychology, complex spatial reasoning goes beyond simple relational judgments, encompassing dynamic world-knowledge reasoning, interactive spatial behavior with environments or agents, logical analysis of three-dimensional structures, and perspective-taking abilities.
+Motivated by these challenges, we introduce OmniSpatial, a comprehensive benchmark designed to capture the breadth and depth of spatial cognition.
+OmniSpatial systematically categorizes spatial reasoning into four core dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—thus providing a principled foundation for developing next-generation spatially- and physically-aware AI systems.
+Table 1: Comparison with other spatial reasoning benchmarks.
+A comparison between OmniSpatial and other existing spatial reasoning benchmarks.
+OmniSpatial avoids template-based annotations, features highly diverse data, and includes a significantly larger number of tasks.
 Dataset Embodied Task Categories Data Domain Data Annotation Data Scale Spatial QAs
-EmbSpatial-Bench [26] 6 Indoor (ScanNet, etc.) Template 2.2K 3.6K
-Space3D-Bench [104] 6 Indoor (Replica) Manual 211 1K
-Visual Spatial [61] 7 MSCOCO Template 10K 10K
-SpatialRGPT-Bench [17] 12 Urban, Indoor, Sim Template 1.4K 1.4K
-BLINK-Spatial [34] 14 MSCOCO Manual 286 286
-What’s up [48] 6 Household, GQA, COCO Template 5K 5K
-Spatial-MM [101] 4 Internet Template 2.3K 2.3K
-RoboSpatial [102] 4 Indoor, tabletop Template 1M 3M
-SpatialVLM [15] 2 WebLi Template 546 546
-SpatialBot-Bench [11] 5 COCO,VG,RTX Manual 200 360
-VSI-Bench [121] 8 Indoor Template 288 5K
-OmniSpatial (Ours) 50 Internet Manual 1387 1.5K
-Table 1: Comparison with other spatial reasoning benchmarks.
-A comparison between OmniSpatial and other existing spatial reasoning benchmarks reveals that, under the inclusion of Embodied Frames Compatibility, OmniSpatial avoids template-based annotations, features highly diverse data, and includes a significantly larger number of tasks.
+EmbSpatial-Bench (Du et al., 2024) 6 Indoor (ScanNet, etc.) Template 2.2K 3.6K
+Space3D-Bench (Szymanska et al., 2024) 6 Indoor (Replica) Manual 211 1K
+Visual Spatial (Liu et al., 2023a) 7 MSCOCO Template 10K 10K
+SpatialRGPT-Bench (Cheng et al., 2024) 12 Urban, Indoor, Sim Template 1.4K 1.4K
+What’s up (Kamath et al., 2023) 6 Household, GQA, COCO Template 5K 5K
+Spatial-MM (Shiri et al., 2024) 4 Internet Template 2.3K 2.3K
+RoboSpatial (Song et al., 2024) 4 Indoor, tabletop Template 1M 3M
+SpatialVLM (Chen et al., 2024) 2 WebLi Template 546 546
+SpatialBot-Bench (Cai et al., 2024) 5 COCO,VG,RTX Manual 200 360
+VSI-Bench (Yang et al., 2024) 8 Indoor Template 288 5K
+OmniSpatial (Ours) 50 Internet Manual 6.5K 8.4K
 The OmniSpatial benchmark includes images or video frames across diverse scenes, resolutions, lighting conditions, and weather patterns, collected from multiple countries across different continents.
@@
 2 Preliminaries: Visual–Spatial Reasoning
-Spatial reasoning constitutes the cognitive bridge between visual perception and action planning in embodied agents.
-We define visual–spatial reasoning as the capacity of an artificial system to infer, predict, and manipulate spatial properties of the world from visual observations.
-Formally, let an RGB observation stream be $\mathbf{I}_{1:T}$ and a task-specific query be $q$.
-A model possesses visual–spatial reasoning if it can learn a mapping
-$f:(\mathbf{I}_{1:T},\,q)\;\longrightarrow\;a,$
-where $a$ belongs to a well-defined action or answer space whose correctness can be verified in the physical or simulated environment.
-This definition excludes non-visual priors so that improvements can be attributed to visual reasoning itself, yet it remains compatible with multi-modal extensions discussed in §1.
+Spatial reasoning constitutes the cognitive bridge between visual perception and geometric understanding.
+We define visual–spatial reasoning as the capacity of an artificial system to infer, predict, and reason spatial properties of the world from visual observations.
+Formally, let an RGB observation stream be $\mathbf{I}_{1:T}$ and a task‑specific query be $q$.
+A model possesses visual–spatial reasoning that learn a mapping:
+$f:(\mathbf{I}_{1:T},\,q)\;\longrightarrow\;a,$ (1)
+where $a$ belongs to a well‑defined action or answer space whose correctness can be verified in the physical or simulated environment.
+This definition excludes non‑visual priors so that improvements can be attributed to visual reasoning itself, yet it remains compatible with multi‑modal extensions discussed in §1.
 2.1 Taxonomy of Visual–Spatial Reasoning
-Our classification stems from two complementary motivations.
-(i) Cognitive-psychology foundations.
-Decades of research on spatial cognition isolate partly independent faculties—e.g., spatial visualization, mental rotation, perspective taking, and spatial updating—that can be psychometrically evaluated in humans [13, 71].
-These constructs offer a principled scaffold for analysing how agents perceive, reason about, and act within space.
-(ii) The need to move beyond basic spatial relations.
-Current spatial benchmarks are nearly saturated on tasks such as left–right discrimination, front–back identification, and object counting [15, 11, 17, 26, 104].
-Real-world embodied tasks, however, require richer reasoning [50, 82, 16, 103, 53] about scene dynamics, multi-step spatial logic, physically grounded interaction, and viewpoint transformation.
-Guided by these two strands, we partition spatial reasoning into four orthogonal yet complementary dimensions—dynamic reasoning, complex spatial logic, spatial interaction, and perspective taking.
-Each dimension aligns with a specific cognitive faculty and targets a distinct class of spatial challenges that is under-represented in prior work, enabling us to construct benchmarks that probe the full spectrum of complex spatial cognition while remaining anchored in established psychological theory.
-Dynamic Reasoning involves predicting motion and temporal changes in spatial environments.
-This capability is crucial for tasks such as robotic control and autonomous navigation, where anticipating object trajectories ensures adaptive decision-making in dynamic contexts.
-Complex Spatial Logic encompasses high-level reasoning about spatial relationships, transformations, and geometric structures.
-It is fundamental to problem-solving in domains such as engineering, design, and robotic manipulation, where understanding spatial configurations and predicting structural changes are essential.
-Spatial Interaction focuses on engagement with spatial environments, where reasoning is guided by environmental constraints and task objectives.
-This includes applications such as path planning, obstacle avoidance, and real-time decision-making based on spatial feedback.
-Perspective Taking refers to the ability to adopt different viewpoints, which is crucial for navigation, social cognition, and spatial awareness.
-It allows for understanding spatial relationships from multiple perspectives, facilitating decision-making in multi-agent environments and adaptive problem-solving.
+Our taxonomy is motivated by two complementary perspectives.
+(i) Cognitive psychology foundations.
+Prior research on spatial cognition highlights partly independent faculties—such as visualization, mental rotation, perspective taking, and spatial updating—that can be systematically assessed in humans (Chabris et al., 2006, Meneghetti et al., 2022).
+These constructs provide a principled scaffold for analyzing how agents perceive, reason about, and act within space.
+(ii) Moving beyond basic spatial relations.
+Existing benchmarks are nearly saturated on simple tasks like left–right discrimination, front–back identification, and object counting (Chen et al., 2024, Cai et al., 2024, Du et al., 2024, Szymanska et al., 2024).
+Yet real-world embodied tasks demand richer reasoning (Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025, Stogiannidis et al., 2025, Lee et al., 2025) about scene dynamics, multi-step logic, physical interaction, and viewpoint transformation.
+Guided by these considerations, we partition visual–spatial reasoning into four complementary dimensions: dynamic reasoning, complex spatial logic, spatial interaction, and perspective taking.
+Each dimension corresponds to a specific cognitive faculty and targets under-explored challenges in prior work, enabling us to probe a broader spectrum of spatial cognition while remaining grounded in psychological theory.
+Dynamic Reasoning concerns inferring motion and temporal change from visual evidence.
+While our benchmark primarily uses static or sparsely sampled frames, such inference is crucial for adaptive decision-making in domains like robotics and navigation.
+Complex Spatial Logic involves higher-order reasoning about relations, transformations, and geometric structures.
+It underpins problem-solving in design, engineering, and manipulation, where anticipating structural or relational changes is essential.
+Spatial Interaction emphasizes reasoning guided by environmental constraints and task goals, covering skills such as path planning, obstacle avoidance, and context-aware action selection.
+Perspective Taking captures the ability to adopt alternative viewpoints, supporting navigation, social cognition, and multi-agent coordination.
+It enables understanding relations from diverse perspectives and fosters flexible problem-solving.
 2.2 Rationale for Classification
-This taxonomy is designed to be both theoretically comprehensive and practically applicable.
-Dynamic reasoning emphasizes motion prediction, which is critical for navigation and interaction.
-Complex spatial logic provides a foundation for advanced cognitive tasks involving transformation and abstraction.
-Spatial interaction highlights real-time engagement with environments, while perspective taking captures cognitive flexibility in spatial reasoning.
-By structuring visual-spatial reasoning in this manner, we establish a framework that enhances the evaluation of spatial cognition models and their applications in artificial intelligence, robotics, and human cognition research.
+This taxonomy balances theoretical comprehensiveness with practical applicability.
+Dynamic reasoning highlights motion inference, complex logic captures abstract transformations, spatial interaction addresses real-time engagement, and perspective taking reflects cognitive flexibility.
+Together, these dimensions provide a framework for evaluating spatial reasoning in AI, robotics, and human cognition.
 Figure 2: Tasks Demonstration of OmniSpatial.
@@
 Note: The questions above are slightly simplified for clarity and conciseness.
+3 OmniSpatial: Comprehensive Spatial Reasoning Benchmark
+3.1 Overview
+We present OmniSpatial, a comprehensive benchmark designed to evaluate vision-language models on spatial reasoning.
+Rather than pursuing sheer data volume, OmniSpatial emphasizes diversity, structure, and rigor.
+It now consists of 8.4K carefully curated question–answer pairs, substantially larger than earlier prototypes, and covers a broad spectrum of scenarios that demand reasoning beyond pattern recognition.
+The dataset integrates heterogeneous sources—web imagery, standardized cognitive tests, driving-exam questions, and prior dataset images such as MME (Liang et al., 2024) and HOI4D (Liu et al., 2022).
+This mixture enriches both realism and complexity: natural images capture everyday environments and architectures; psychology-inspired tasks introduce scientifically grounded challenges; driving exams provide safety-critical dynamic reasoning; and embodied datasets contribute varied resolutions, viewpoints, and human–object interactions.
+Tasks are organized into 4 major categories and 50 fine-grained subtypes as shown in Figs.˜6 and 2, spanning from basic perspective-taking to dynamic motion prediction and spatial interaction in cluttered scenes.
+Each item is manually designed and reviewed through multi-round annotation, ensuring accuracy, consistency, and minimal ambiguity.
+This careful curation yields a benchmark that not only broadens the coverage of spatial reasoning but also establishes a reliable ground for evaluating and advancing future multimodal intelligence.
 Figure 3: Data Construction of OmniSpatial.The pipeline collects images from multiple sources and ensures their quality and relevance through manual selection.
@@
 This process supports the effective training of VLMs in spatial reasoning tasks.
-3 OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark
-3.1 Overview
-We introduce OmniSpatial, a benchmark for the quantitative evaluation of VLMs in visual-spatial reasoning.
-Unlike large-scale benchmarks focused on a few tasks, OmniSpatial prioritizes task diversity and structured categorization over sheer dataset size.
-It consists of 1533 high-quality question-answer pairs, designed to challenge models beyond simple pattern matching or statistical learning, ensuring that spatial reasoning capabilities are tested rigorously, even with limited data.
-Our dataset is curated from web searches, standardized tests, driving exam questions, and existing datasets including MME [57], HOI4D [67], etc.This diverse sourcing enhances the realism, complexity, and cross-domain generalization of the tasks.
-Web-sourced images cover natural environments, architecture, and daily life, adding visual complexity.
-Standardized tests from psychology and cognitive science provide scientifically rigorous spatial reasoning challenges.
-Driving exam questions introduce real-world dynamic interactions, such as road rule comprehension and motion prediction.
-Additionally, datasets like HOI4D contribute multi-resolution, varying illumination, and viewpoint diversity, as well as embodied intelligence tasks involving human-object interactions.
-To ensure comprehensive evaluation, we categorize tasks into 4 major spatial reasoning types, further divided into 50 fine-grained task categories, as shown in Figs.
-6 and 2.
-For example, perspective-taking tasks range from simple spatial judgments (“Is the opponent’s weapon to the left or right of the blue athlete?”) to complex motion prediction (“The blue athlete is blocking an incoming attack from the red athlete—what will happen next?”).
-Spatial interaction tasks include both static collision assessments (“Is the vehicle too close to the one in front?”) and dynamic environment reasoning (“A car door ahead is opening—a pedestrian exiting?”).
-This multi-level task structure ensures models are tested on diverse and progressively complex spatial reasoning challenges.
-We emphasize precise image-task alignment in both dataset selection and annotation.
-Each question-answer pair is manually curated and reviewed in multiple rounds to ensure accuracy, consistency, and minimal ambiguity.
-Unlike automated annotation, manual refinement guarantees a high correlation between visual scenes and spatial reasoning tasks, strengthening OmniSpatial’s reliability as a benchmark for future research.
 3.2 Benchmark Construction
-This section describes how we collect data and annotations for OmniSpatial, as illustrated in Fig. 3.
 3.2.1 Data Collection
-As outlined in Section 2, our study defines 4 spatial categories with corresponding task types.
-To ensure a diverse and information-rich dataset, we design targeted search strategies for each task type, optimizing data collection for relevance, diversity, and complexity.
-Web Images Web images constitute a major portion of our dataset, particularly for Perspective Taking, Dynamic Reasoning, and Spatial Interaction tasks.
-To ensure relevance and diversity, we formulate task-specific search terms reflecting distinct spatial scenarios, such as “indoor layout” and “furniture arrangement” for spatial relationship tasks.
-To minimize AI-generated image interference, we append filters like “-ai” and “-generated” to search queries.
-Using Google’s Custom Search JSON API, Web Robotic Process Automation (Web RPA), and manual searches, we retrieve thousands of images efficiently.
-A rigorous screening process is followed to eliminate irrelevant, low-resolution, or overly compressed images.
-We also removed images lacking spatial complexity, such as isolated static objects in Spatial Interaction tasks.
-The final dataset balances visual complexity and realism, enhancing its generalization capability across multiple tasks and scenarios while ensuring high-quality training data for VLMs.
-All the images are under MIT or CC-BY 4.0 licence.
-Exam-Based Test Questions  Designed to assess complex spatial reasoning, these questions emphasize abstract spatial logic, including 3D transformations, rotations, and perspective shifts.
-We collect a wide range of publicly available spatial cognition tests via web scraping and manual curation.
-Noting variations in task focus, difficulty, and reasoning style, we categorized the questions to maintain a balanced representation.
-To prevent overrepresentation of any single reasoning type, we curate the dataset to ensure diverse question formats.
-Redundant, overly similar, or knowledge-intensive questions are removed, prioritizing those that purely test spatial reasoning.
-This refinement process significantly enhances the dataset’s quality and challenge level, providing a robust evaluation framework for VLMs.
-Driving Test Questions  These questions evaluate spatial interaction in dynamic environments, focusing on distance estimation, right-of-way rules, and hazard detection.
-We source them from three primary channels: (i) Image-based multiple-choice questions from driving test websites in at least three countries, capturing diverse traffic rules and scenarios.
-(ii) Online question banks covering standardized spatial judgment tasks such as turning, lane changes, and parking decisions.
-(iii) Interactive driving videos from U.S.
-test resources, where we extract frames, annotate bounding boxes, and formulate contextual questions (“Which bounding box highlights a potential traffic risk?”).
-This multi-source approach ensures a realistic and challenging dataset, enhancing VLM adaptability in traffic-related spatial reasoning.
-Existing Datasets  We integrate two key datasets: MME [57] and HOI4D [67].
+As described in Section˜2, we define four spatial categories and corresponding task types.
+To build a diverse and information-rich dataset, we design targeted search strategies for each type, optimizing for relevance, diversity, and complexity.
+Web Images.
+Web images form a major part of our data, especially for perspective-taking, dynamic reasoning, and spatial interaction.
+We use task-specific search terms (e.g., “indoor layout,” “furniture arrangement”) and append filters (“-ai,” “-generated”) to reduce synthetic content.
+Images are retrieved via Google’s Custom Search JSON API, Web RPA, and manual search, followed by strict filtering to remove irrelevant, low-resolution, or spatially trivial cases (e.g., isolated static objects).
+The resulting set balances realism and complexity, ensuring broad task coverage.
+All images are under MIT or CC-BY 4.0 license.
+Exam-Based Test Questions.
+To capture abstract spatial logic, such as 3D transformations, rotations, and perspective shifts, we collect public spatial cognition tests through web scraping and manual curation.
+We categorize questions by focus and difficulty to maintain balance, removing redundant or knowledge-heavy items in favor of those targeting pure spatial reasoning.
+This refinement increases challenge diversity and improves benchmark quality.
+Driving Test Questions.
+To evaluate reasoning in dynamic environments, we source tasks from three channels: (i) image-based multiple-choice questions from driving exam websites across at least three countries, (ii) online banks of standardized tasks like turning, lane changing, and parking, and (iii) interactive U.S.
+driving test videos, from which we extract frames, annotate bounding boxes, and design contextual queries (e.g., “Which bounding box indicates a potential traffic hazard?”).
+This combination yields realistic and challenging traffic scenarios, enhancing VLM adaptability to safety-critical reasoning tasks.
+Existing Dataset Images  We integrate two key data sources: MME (Liang et al., 2024) and HOI4D (Liu et al., 2022).
 MME provides RGB-D data, allowing depth-based spatial inference.
-We leverage its depth information for physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning.
+We leverage its depth information and manually propose physics-based questions such as “If a red car passes me in 5 seconds, what speed should I maintain?” This ensures realistic distance and motion-based reasoning.
 HOI4D contains extensive human-object interaction videos.
 We extract sequential frames to create motion prediction tasks, such as “Where will the hand holding the kettle move next?” By incorporating these datasets, we introduce real-world motion and interaction complexities, further strengthening VLMs’ dynamic reasoning capabilities.
+Furthermore, to extend our dataset to model training, we partition the dataset into a 1.5K test set and a 6.9K training set.
+The test set is entirely human-annotated and further integrates samples from several existing datasets.
+including SpatialViz  (Wang et al., 2025), PhysBench (Chow et al., 2025), ViewSpatial (Li et al., 2025), and DrivingVQA (Corbière et al., 2025), which substantially enhance the diversity of our dataset.
 Figure 4: PointGraph: Enhance Spatial Reasoning through Additional Scene Graphs.
@@
 3.2.2 Question-Answer Annotation
-We employ multiple-choice questions, including binary (true/false) and four-option formats, ensuring standardized evaluation and minimizing annotation errors.
-While adhering to structured formats, we prioritize natural and contextually rich language to prevent overly templated expressions.
-Instead of rigid, mathematical phrasing (e.g., “Is [A] object on the right of [B] object?”), We use conversational styles more aligned with real-world reasoning (e.g., “If you’re entering the classroom, on which side are the students?”).
-This approach encourages models to rely on contextual and relative spatial references rather than predefined templates.
-To maintain clarity, we conduct multiple rounds of validation, resolving ambiguities in referents and ensuring unique answers.
-In particular, we ensure the quality of the data by means of cross-validation among different annotators.
-We employ six human annotators for data labeling.
-All annotators receive comprehensive training to ensure a thorough understanding of the spatial reasoning tasks, enabling them to provide unbiased and discriminative assessments.
-All the annotators are from the authors of this work.
+We design multiple-choice questions, including binary (true/false) and four-option formats, to enable standardized evaluation while minimizing annotation bias.
+To ensure naturalness, questions are phrased in conversational and context-rich styles (e.g., “If you are entering the classroom, on which side are the students?”) rather than rigid templates (e.g., “Is [A] on the right of [B]?”).
+This encourages models to rely on contextual and relational reasoning rather than memorized patterns.
+To guarantee clarity and answer uniqueness, we perform multiple rounds of validation and resolve ambiguous references through cross-checking among annotators.
+Six trained annotators—initially the authors—were involved in labeling.
+Each annotator underwent task-specific training, and annotations were cross-validated to reduce subjectivity.
+The inter-annotator agreement reached Krippendorff’s $\alpha=0.84$, indicating a high level of consistency.
 3.3 Improving Visual Spatial Reasoning Abilities
-Given the diversity of spatial reasoning tasks in OmniSpatial, we further explore methods to enhance VLMs’ spatial reasoning capabilities.
-3.3.1 PointGraph: Enhancing Spatial Reasoning via Point Relationships
-We attempt to employ existing expert models, such as the Segment Anything Model [49], to partition images into distinct regions.
-This process extracts pixel clusters corresponding to multiple objects in the image, allowing models to estimate object centers and perform subsequent spatial reasoning tasks more effectively.
-As shown in Fig.
-5, We attempt to enhance the spatial relationship comprehension of VLMs by explicitly incorporating additional spatial information cues into the models.
-3.3.2 SpatialCoT: Enhancing Spatial Imagination via Novel View Synthesis
-We observe that VLMs are largely constrained by textual reasoning paradigms, whereas humans inherently possess robust spatial imagination, commonly known as “mental imagery”.
-To some extent, this capability of spatial imagination corresponds to novel view synthesis of objects or scenes.
-Fortunately, recent advances in 3D generative models [100, 118, 87, 66] can perform such tasks effectively.
-In our work, we employ InstantMesh [118] to synthesize novel views from input images, providing additional visual cues to enhance spatial imagination.
-Method Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+3.3.1 PointGraph: Explicit Modeling of Object Relationships
+To strengthen models’ understanding of spatial relations, we introduce PointGraph, which constructs structured representations of object relationships within an image.
+Concretely, we employ open-vocabulary grounding models such as Florence-2 to localize multiple objects and extract their centers and bounding boxes.
+These detections are then assembled into a JSON-style scene graph encoding object identities and relative positions.
+By concatenating this structured spatial description with the original query, VLMs are provided with explicit geometric cues that facilitate more accurate reasoning about distances, directions, and configurations (see Fig.˜5).
+3.3.2 SpatialCoT: Stimulating Spatial Imagination via Novel Views
+Human spatial reasoning often relies on mental imagery—the ability to imagine how a scene looks from different viewpoints.
+Inspired by this, we design SpatialCoT, which augments visual inputs with 3D novel views to enrich spatial imagination (Shi et al., 2023, Qi et al., 2023b, Liu et al., 2024d).
+Specifically, we adopt InstantMesh (Xu et al., 2024a) to synthesize six additional perspectives for each input image and compose them into a multi-view collage.
+This collage is then fed, along with the question, into the VLM as part of a chain-of-thought prompting pipeline.
+The additional perspectives provide strong geometric priors, helping models disambiguate occlusions, perspective-taking, and other view-dependent reasoning tasks.
+Table 2: Evaluation on OmniSpatial-test.
+All models were tested 5 times and averaged to reduce randomness.
+Dark green indicates the best result and light green indicates the second-best result within the group.
+Gemini-2.5-Pro (Reid et al., 2024), InternVL3-78B (Zhu et al., 2025), and SoFar (Qi et al., 2025) achieve the best performance within their respective groups.
+Method
+Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
 Manipulate
@@
 Random Choice 24.98 - 24.86 26.30 25.88 23.43 27.27 21.44 24.77 22.55 24.84 25.78
-GPT-3.5-turbo [98] 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
-GPT-4-turbo [77] 34.06 - 42.97 37.40 41.18 28.95 40.00 22.27 26.32 31.37 33.99 35.42
+GPT-3.5-turbo (Roumeliotis & Tselikas, 2023) 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
+GPT-4-turbo (OpenAI, 2023) 34.06 - 42.97 37.40 41.18 28.95 40.00 22.27 26.32 31.37 33.99 35.42
 Proprietary Models
-GPT-4o-mini-2024-07-18 [45] 42.64 8 55.95 50.29 54.59 43.43 44.91 22.47 29.42 61.57 36.76 34.22
-GPT-4o-2024-11-20 [45] 47.81 5 65.54 57.23 56.47 52.38 54.09 26.29 25.48 75.98 39.49 39.76
-GPT-4.1-nano-2025-04-14 [78] 42.62 9 50.90 53.85 54.90 40.95 42.42 24.40 30.11 53.59 37.23 33.73
-GPT-4.1-mini-2025-04-14 [78] 48.87 4 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
-GPT-4.1-2025-04-14 [78] 51.78 2 66.22 64.74 60.00 65.33 60.18 31.75 30.06 70.98 40.64 39.04
-Claude-3-5-sonnet-20241022 [3] 46.86 6 54.05 54.57 58.12 68.38 53.09 26.60 31.74 70.00 34.79 39.52
-Claude-3-7-sonnet-20250219 [3] 47.53 5 57.57 55.95 56.71 63.81 59.09 29.48 28.39 72.16 36.06 36.63
-Gemini-2.0-flash-lite-02-05 [2] 44.03 8 59.19 46.71 60.24 49.52 53.27 21.65 31.23 66.47 36.81 38.80
-Gemini-2.0-flash-exp [2] 48.40 2 61.89 56.01 51.76 63.43 59.09 20.82 33.81 72.75 39.20 39.28
-Gemini-2.5-flash-preview-05-20 [2] 52.12 1 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
+GPT-4o-mini-2024-07-18 (Hurst et al., 2024) 42.64 8 55.95 50.29 54.59 43.43 44.91 22.47 29.42 61.57 36.76 34.22
+GPT-4o-2024-11-20 (Hurst et al., 2024) 47.81 5 65.54 57.23 56.47 52.38 54.09 26.29 25.48 75.98 39.49 39.76
+GPT-4.1-nano-2025-04-14 (OpenAI, 2025a) 42.62 9 50.90 53.85 54.90 40.95 42.42 24.40 30.11 53.59 37.23 33.73
+GPT-4.1-mini-2025-04-14 (OpenAI, 2025a) 48.87 4 64.32 56.53 59.06 60.19 56.36 29.28 30.19 72.55 39.57 39.28
+GPT-4.1-2025-04-14 (OpenAI, 2025a) 51.78 2 66.22 64.74 60.00 65.33 60.18 31.75 30.06 70.98 40.64 39.04
+Claude-3-5-sonnet-20241022 (Anthropic, 2024) 46.86 6 54.05 54.57 58.12 68.38 53.09 26.60 31.74 70.00 34.79 39.52
+Claude-3-7-sonnet-20250219 (Anthropic, 2024) 47.53 5 57.57 55.95 56.71 63.81 59.09 29.48 28.39 72.16 36.06 36.63
+Gemini-2.0-flash-lite-02-05 (Anil et al., 2023) 44.03 8 59.19 46.71 60.24 49.52 53.27 21.65 31.23 66.47 36.81 38.80
+Gemini-2.0-flash-exp (Anil et al., 2023) 48.40 2 61.89 56.01 51.76 63.43 59.09 20.82 33.81 72.75 39.20 39.28
+Gemini-2.5-flash-preview-05-20 (Anil et al., 2023) 52.12 1 67.57 62.72 68.24 73.33 60.91 38.14 34.19 75.49 35.90 33.73
 Reasoning Models
-o1-2024-12-17 [46] 50.36 6 71.62 60.98 57.65 63.81 60.00 39.18 27.10 71.57 38.03 36.14
-o4-mini-2025-04-16 [79] 52.77 3 72.97 59.83 60.00 73.33 61.82 34.02 36.77 73.53 40.69 40.96
-o3-2025-04-16 [79] 56.33 1 71.89 66.18 61.18 68.57 65.45 40.21 29.68 77.06 48.40 48.19
-Claude-3-7-sonnet-20250219-thinking [3] 48.62 7 57.21 59.73 53.73 67.94 57.27 30.24 28.17 68.63 37.94 36.95
-Gemini-2.5-flash-05-20-thinking [2] 53.16 3 70.27 64.74 61.18 72.38 58.18 35.05 36.13 74.12 40.96 32.53
-Gemini-2.5-pro-preview-05-06 [2] 55.19 2 67.57 71.39 62.35 75.24 64.55 43.30 34.84 74.51 38.03 37.35
+o1-2024-12-17 (Jaech et al., 2024) 50.36 6 71.62 60.98 57.65 63.81 60.00 39.18 27.10 71.57 38.03 36.14
+o4-mini-2025-04-16 (OpenAI, 2025b) 52.77 3 72.97 59.83 60.00 73.33 61.82 34.02 36.77 73.53 40.69 40.96
+o3-2025-04-16 (OpenAI, 2025b) 56.33 1 71.89 66.18 61.18 68.57 65.45 40.21 29.68 77.06 48.40 48.19
+Claude-3-7-sonnet-20250219-thinking (Anthropic, 2024) 48.62 7 57.21 59.73 53.73 67.94 57.27 30.24 28.17 68.63 37.94 36.95
+Gemini-2.5-flash-05-20-thinking (Anil et al., 2023) 53.16 3 70.27 64.74 61.18 72.38 58.18 35.05 36.13 74.12 40.96 32.53
+Gemini-2.5-pro-preview-05-06 (Anil et al., 2023) 55.19 2 67.57 71.39 62.35 75.24 64.55 43.30 34.84 74.51 38.03 37.35
 Open-source Models
-LLavA-1.5-vicuna-7B [64] 34.97 15 54.46 31.23 35.29 36.19 33.94 29.01 24.18 55.60 34.66 36.14
-LLaVA-onevision-qwen2-7B [54] 35.68 14 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
-LLaVA-onevision-qwen2-72B [54] 45.66 6 62.16 50.29 54.12 60.95 56.36 22.68 25.81 76.47 37.23 33.73
-Gemma-3-4B [47] 39.79 11 41.89 49.71 56.47 27.62 36.36 23.71 24.52 59.80 36.17 38.55
-Gemma-3-12B [47] 43.71 8 54.05 54.91 54.12 47.62 45.45 16.49 30.32 63.73 36.70 33.73
-Gemma-3-27B [47] 44.75 7 56.76 55.78 57.65 50.48 52.73 27.84 29.03 64.71 33.51 32.53
-InternVL3-2B [130] 37.98 13 50.00 40.58 43.29 40.00 40.55 21.86 28.52 55.49 35.11 33.01
-InternVL3-8B [130] 41.60 9 52.43 40.87 48.94 51.05 44.77 24.95 28.63 64.20 38.62 40.96
-InternVL3-14B [130] 45.94 5 54.32 60.17 50.35 51.81 51.45 28.04 28.26 68.04 35.37 34.46
-InternVL3-38B [130] 48.48 2 63.42 63.58 54.59 58.29 50.55 29.90 28.52 72.16 36.76 33.49
-InternVL3-78B [130] 49.33 1 63.78 63.12 56.24 59.24 51.45 27.63 30.19 74.51 38.46 35.90
-Qwen-VL2.5-3B [115] 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
-Qwen-VL2.5-7B [115] 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
-Qwen-VL2.5-32B [115] 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
-Qwen-VL2.5-72B [115] 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
+LLavA-1.5-vicuna-7B (Liu et al., 2024c) 34.97 15 54.46 31.23 35.29 36.19 33.94 29.01 24.18 55.60 34.66 36.14
+LLaVA-onevision-qwen2-7B (Li et al., 2024a) 35.68 14 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
+LLaVA-onevision-qwen2-72B (Li et al., 2024a) 45.66 6 62.16 50.29 54.12 60.95 56.36 22.68 25.81 76.47 37.23 33.73
+Gemma-3-4B (Kamath et al., 2025) 39.79 11 41.89 49.71 56.47 27.62 36.36 23.71 24.52 59.80 36.17 38.55
+Gemma-3-12B (Kamath et al., 2025) 43.71 8 54.05 54.91 54.12 47.62 45.45 16.49 30.32 63.73 36.70 33.73
+Gemma-3-27B (Kamath et al., 2025) 44.75 7 56.76 55.78 57.65 50.48 52.73 27.84 29.03 64.71 33.51 32.53
+InternVL3-2B (Zhu et al., 2025) 37.98 13 50.00 40.58 43.29 40.00 40.55 21.86 28.52 55.49 35.11 33.01
+InternVL3-8B (Zhu et al., 2025) 41.60 9 52.43 40.87 48.94 51.05 44.77 24.95 28.63 64.20 38.62 40.96
+InternVL3-14B (Zhu et al., 2025) 45.94 5 54.32 60.17 50.35 51.81 51.45 28.04 28.26 68.04 35.37 34.46
+InternVL3-38B (Zhu et al., 2025) 48.48 2 63.42 63.58 54.59 58.29 50.55 29.90 28.52 72.16 36.76 33.49
+InternVL3-78B (Zhu et al., 2025) 49.33 1 63.78 63.12 56.24 59.24 51.45 27.63 30.19 74.51 38.46 35.90
+Qwen-VL2.5-3B (Wang et al., 2024b) 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
+Qwen-VL2.5-7B (Wang et al., 2024b) 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
+Qwen-VL2.5-32B (Wang et al., 2024b) 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
+Qwen-VL2.5-72B (Wang et al., 2024b) 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
 Specialized Spatial Reasoning Models
-SpaceMantis-13B [15] 36.36 6 47.03 36.59 40.94 34.86 33.09 22.27 24.39 49.22 38.25 39.28
-SpaceQwen2.5-VL-3B [15] 40.25 3 58.11 39.88 41.18 40.95 40.91 29.90 25.81 63.73 38.83 39.76
-SpaceThinker-Qwen2.5VL-3B [15] 40.42 2 47.84 53.06 43.29 35.43 38.73 24.33 28.00 58.04 35.11 31.08
-SpatialBot-3B [11] 35.68 6 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
-RoboPoint-vicuna-v1.5-7B-lora [123] 35.85 6 57.03 28.61 34.82 37.33 40.55 29.90 22.71 50.20 38.72 40.96
-RoboPoint-vicuna-v1.5-13B [123] 34.60 5 55.68 28.15 42.82 32.19 32.55 24.12 27.74 49.02 37.66 33.49
-SoFar-Qwen2.5VL-3B [89] 45.14 1 56.49 51.16 54.12 53.14 52.73 31.75 22.88 71.60 36.56 41.69
+SpaceMantis-13B (Chen et al., 2024) 36.36 6 47.03 36.59 40.94 34.86 33.09 22.27 24.39 49.22 38.25 39.28
+SpaceQwen2.5-VL-3B (Chen et al., 2024) 40.25 3 58.11 39.88 41.18 40.95 40.91 29.90 25.81 63.73 38.83 39.76
+SpaceThinker-Qwen2.5VL-3B (Chen et al., 2024) 40.42 2 47.84 53.06 43.29 35.43 38.73 24.33 28.00 58.04 35.11 31.08
+SpatialBot-3B (Cai et al., 2024) 35.68 6 43.24 38.15 32.94 29.52 41.82 28.87 22.58 47.06 36.17 37.35
+RoboPoint-vicuna-v1.5-7B-lora (Yuan et al., 2024) 35.85 6 57.03 28.61 34.82 37.33 40.55 29.90 22.71 50.20 38.72 40.96
+RoboPoint-vicuna-v1.5-13B (Yuan et al., 2024) 34.60 5 55.68 28.15 42.82 32.19 32.55 24.12 27.74 49.02 37.66 33.49
+SoFar-Qwen2.5VL-3B (Qi et al., 2025) 45.14 1 56.49 51.16 54.12 53.14 52.73 31.75 22.88 71.60 36.56 41.69
 Human Evaluation
-Human 92.63 - 96.53 97.30 92.94 97.14 94.55 91.30 87.63 99.02 95.74 93.98
-Table 2: Evaluation on OmniSpatial.
-All models were tested 5 times and averaged to reduce randomness.
-Dark green indicates the best result within the group, and light green indicates the second-best result within the group.
-Gemini-2.5-Pro [97], InternVL3-78B [130], and SoFar [89] achieve the best performance within their respective groups.
-Among the various spatial intelligence categories, Complex Logic is the most challenging.
+Human 92.63 - 94.62 96.07 91.38 95.11 92.15 89.02 85.90 98.53 94.30 90.26
 4 Experiments
 4.1 Evaluation Setup
-To assess the spatial reasoning capabilities of existing VLMs, we conduct comprehensive experiments on OmniSpatial.
-This evaluation compares both proprietary and open-source models across various spatial tasks, using standardized benchmarks to ensure fairness and reproducibility.
-Benchmark Models  We evaluate a range of state-of-the-art vision-language models, including:
-  • Proprietary Models: These include GPT-4o series [45], GPT-4.1 series, Claude series [3], and Gemini series [2], which are state-of-the-art multimodal models accessed via an API under zero-shot settings with standardized prompts.
-All the system prompts can be found at  Appendix D.
-  • Reasoning Models: We refer to current models employing internal long CoT as reasoning models.
-These models typically leverage reinforcement learning to enhance their reasoning capabilities, including models such as the o1, o3, o4-mini [46, 79], and Gemini-2.5 series [2, 97].
-Due to the explicit nature of CoT, we utilize an additional text-based evaluation model [98] to determine the correctness of responses, similar to LLM-as-a-Judge [129].
-  • Open Source Models: Our open-source set comprises Qwen-VL series [6], InternVL series [130], Gemma series [47] and LLaVA-Onevision series [63].
-These models are currently the most advanced open-source VLMs.
-All models are deployed locally and use standardized prompts.
-  • Specialized Spatial Models: These models, including SpatialVLM [15], RoboPoint [123], SpatialBot [11], and SoFar [89], are designed specifically for spatial reasoning tasks.
-SpatialVLM integrates metric spatial information through 3D data synthesis and scene segmentation.
-RoboPoint and SpatialBot emphasize the role of point affordance in spatial reasoning.
+To systematically assess spatial reasoning, we evaluate models on the OmniSpatial benchmark under a unified protocol.
+Our evaluation covers both proprietary and open-source systems, spanning general-purpose VLMs, reasoning-oriented LLMs, and spatially specialized models, thereby ensuring broad coverage and fair comparison.
+We consider four groups of state-of-the-art models:
+  • Proprietary Models.
+These include the GPT-4o and GPT-4.1 families (Hurst et al., 2024), the Claude series (Anthropic, 2024), and Gemini series (Anil et al., 2023).
+All are accessed through APIs under zero-shot settings with standardized system prompts (see Appendix˜D).
+  • Reasoning Models.
+We categorize models that explicitly employ long chain-of-thought reasoning, often enhanced through reinforcement learning, such as o1, o3, o4-mini (Jaech et al., 2024, OpenAI, 2025b), and Gemini-2.5 (Anil et al., 2023, Reid et al., 2024).
+Because their outputs are less amenable to strict parsing, we additionally rely on an automatic judge to compare answers against ground truth, following the LLM-as-a-Judge paradigm (Zheng et al., 2023).
+  • Open-Source Models.
+This set includes Qwen-VL (Bai et al., 2023), InternVL (Zhu et al., 2025), Gemma (Kamath et al., 2025), and LLaVA-OneVision (Liu et al., 2023b).
+These are locally deployed with standardized prompts to ensure reproducibility.
+  • Specialized Spatial Models.
+We further benchmark models designed specifically for spatial reasoning, including SpatialVLM (Chen et al., 2024), RoboPoint (Yuan et al., 2024), SpatialBot (Cai et al., 2024), and SoFar (Qi et al., 2025).
+These systems incorporate explicit spatial signals such as metric 3D information, point affordances, or semantic orientation to improve reasoning.
 4.2 Evaluation Metrics
-We evaluate the spatial reasoning capabilities of models by measuring their accuracy on multiple-choice questions.
-For standard open-source and closed-source models, we experiment with four approaches: enforcing direct output result, regular expression parsing, JSON format parsing and LLM-as-a-Judge [129].
-For Reinforcement Learning-based reasoning models, due to the inherent challenges in enforcing structured output, we employ GPT-4.1-mini to assess the correctness of the model’s responses by comparing them with the ground truth.
-More detailed ablation studies are discussed in the Section C.1.
-Method Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+We measure accuracy on multiple-choice questions.
+For standard proprietary and open-source models, we test four output protocols: direct answer, regular-expression parsing, JSON parsing, and LLM-as-a-Judge (Zheng et al., 2023).
+For reasoning-oriented models with unstructured CoT outputs, correctness is assessed by GPT-4.1-mini against ground truth.
+Ablations of these evaluation strategies are provided in Section˜C.1.
+Table 3: Camparison of textual Chain-of-Thought and PointGraph on OmniSpatial-test.
+Method
+Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
 Manipulate
@@
 (w/ PointGraph) 44.36 +2.91 55.68 55.20 48.94 52.19 52.36 29.90 25.55 66.08 35.11 31.08
-Table 3: Camparison of textual Chain-of-Thought and PointGraph on OmniSpatial.
 4.3 Main Results
-Overall Model Performance  As illustrated in Table 2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 [79] and Gemini-2.5-pro [2], achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and require a lot of inference time and tokens.
-(ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B [130] and Qwen-VL2.5-72B [115] achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp.
+Overall Model Performance  As illustrated in Table˜2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 (OpenAI, 2025b) and Gemini-2.5-pro (Anil et al., 2023), achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and they require a lot of inference time and tokens.
+(ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B (Zhu et al., 2025) and Qwen-VL2.5-72B (Wang et al., 2024b) achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp.
 (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
-Category-wise Analysis  Breaking down the results, we observe notable performance differences across spatial reasoning categories: (i) Leveraging their extensive world knowledge and local understanding capabilities, proprietary models have demonstrated strong performance in Dynamic Reasoning and Spatial Interaction, indicating that reasoning models possess high proficiency in temporal understanding, spatial relationship analysis, and map-based comprehension.
-(ii) For Pattern Geometric and Recognition Reasoning, which involve spatial imagination in planar geometry, even reasoning models designed for extended thinking can only achieve an accuracy of around 30% to 40%, slightly surpassing the random baseline.
-(iii) Current models exhibit limited Perspective Taking abilities, predominantly analyzing scenarios from an ego-centric viewpoint while struggling to imagine perspectives from others’ viewpoints.
+Category-wise Analysis  We observe notable performance differences across spatial reasoning categories: (i) Leveraging their extensive world knowledge and local understanding capabilities, proprietary models have demonstrated strong performance in Dynamic Reasoning and Spatial Interaction, indicating that reasoning models possess high proficiency in temporal understanding, spatial relationship analysis, and map-based comprehension.
+(ii) For Pattern Geometric Reasoning, which involves spatial imagination in planar geometry, even reasoning models designed for extended thinking can only achieve an accuracy of around 30% to 40%, slightly surpassing the random baseline.
+(iii) Current models exhibit limited perspective-taking abilities, predominantly analyzing scenarios from an ego-centric viewpoint while struggling to imagine perspectives from others’ viewpoints.
+Table 4: Performance of Spatial CoT on OmniSpatial Perspective-Taking track.
 Method Avg. Improve
@@
 GPT-4.1-mini
-(w/ Zero-shot CoT) 45.56 - 69.41 40.11 40.96
+(w/ Zero‑shot CoT) 45.56 - 69.41 40.11 40.96
 (w/ Spatial CoT) 47.58 +2.02 69.43 42.37 44.34
-Qwen-VL2.5-3B
-(w/ Zero-shot CoT) 40.89 - 59.61 36.54 37.59
+Qwen‑VL2.5‑3B
+(w/ Zero‑shot CoT) 40.89 - 59.61 36.54 37.59
 (w/ Spatial CoT) 42.90 +2.01 60.80 39.25 37.44
-Table 4: Performance of Spatial CoT on OmniSpatial Perspective-Taking track.
 Impact of PointGraph & Spatial CoT  To test whether structured segmentation improves performance, we apply PointGraph as a preprocessing step for GPT-4.1, Gemini-2.5-flash and Qwen-VL2.5-7B.
-Results in Table 3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement.
-Fig.
-5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track.
+Results in Table˜3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement.
+Fig.˜5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track.
 Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.
+4.4 Training Exploration
+Supervised Training on OmniSpatial-train.
+We further study whether OmniSpatial-train can effectively teach spatial skills instead of merely fitting templates.
+Starting from a strong open-source baseline, supervised fine-tuning on 6.9K samples yields a substantial +7.82 point average gain over zero-shot, with consistent improvements across dynamic, interaction, and perspective-taking oriented tracks.
+In contrast, training on a much larger 200K template-style corpus that follows the construction process of VSI-Bench (Yang et al., 2024) brings only a marginal +1.29 average gain, underscoring the value of diverse, manually curated spatial tasks over synthetic templates.
+Table 5: Training Exploration on OmniSpatial-test.
+Method Avg. Improve
+Manipulate
+Motion
+Analysis
+Traffic
+Analysis
+Locate
+Geospatial
+Strategy
+Pattern
+Recognition
+Geometric
+Reasoning
+Ego
+Centric
+Allo
+Centric
+Hypothetical
+Qwen-VL2.5-3B - - - - - - - - - - - -
+Zero-shot 40.30 - 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
++ OmniSpatial-train (6.9K) 48.12 +7.82 59.19 51.16 50.12 51.81 45.45 34.02 31.74 68.63 36.76 35.90
++ Template corpus (200K) 41.59 +1.29 50.90 46.71 40.94 45.33 40.55 28.87 24.39 72.55 35.11 33.01
+Generalization to VSI-Bench.
+To examine cross-benchmark generalization, we adopt the SpaceR-7B pipeline and compare supervised training with and without OmniSpatial.
+Adding OmniSpatial improves the overall score on VSI-Bench from 41.68 to 43.68.
+Notably, it boosts categories requiring ordering, counting, and metric/room-size reasoning (e.g., appearance_order, obj_counting, obj/room_size).
+These results indicate that OmniSpatial provides complementary supervision that transfers to external spatial tasks rather than overfitting to in-benchmark patterns.
+Table 6: Generalization on VSI-Bench. Training with OmniSpatial yields consistent gains.
+Method overall
+Appearance
+Order
+Obj Abs
+Distance
+Obj
+Counting
+Obj Rel
+Distance
+Obj
+Size
+Room
+Size
+Route
+Planning
+Obj Rel
+Direction
+Qwen-VL2.5-3B - - - - - - - - -
+Zero-shot 34.06 34.63 11.39 40.18 36.20 46.11 38.58 31.96 36.92
++ SpaceR-7B 41.68 46.60 19.63 55.27 34.23 58.32 35.17 34.54 42.84
++ OmniSpatial 43.68 58.25 15.13 57.36 34.37 60.99 41.11 34.02 44.16
 5 Related Works
 5.1 Benchmarking Spatial Reasoning
-The development of standardized benchmarks [55, 122, 65, 124, 44, 51] has been critical in assessing the performance of spatial reasoning models.
-Various studies have introduced innovative benchmarking methodologies [104, 15, 17, 11, 102, 89, 96] to advance spatial reasoning evaluation.
-For instance, Spatial VQA [26] was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning.
-SpatialRGPT-Bench [17] further introduced depth-aware reasoning, utilizing large language models to analyze spatial relationships, such as understanding relative directions and distances between specified regions in 3D scenes [104].
-SpatialBench [11] categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks.
-RoboSpatial [102] proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives.
-VSI-Bench [121] combined video data [14, 65, 33, 56, 32] with cognitive maps [72, 4] to simulate human-like spatial cognition and optimize reasoning in dynamic environments.
-Recently, SoFar [89] introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
+Various studies have introduced innovative benchmarking methodologies (Szymanska et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Qi et al., 2025, Ray et al., 2024) to advance spatial reasoning evaluation.
+Spatial VQA (Du et al., 2024) was among the first to incorporate spatial information into vision-language models, enabling fundamental spatial relationship reasoning.
+SpatialBot (Cai et al., 2024) categorized spatial reasoning into various hierarchical levels, extending its applicability to robotic manipulation tasks.
+RoboSpatial (Song et al., 2024) proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives.
+VSI-Bench (Yang et al., 2024) combined video data (Chandrasegaran et al., 2024, Liu et al., 2023c, Fu et al., 2024a, Li et al., 2024b, Fang et al., 2024b) with cognitive maps (Momennejad et al., 2023, Apostolopoulos & Groumpos, 2023) to simulate human-like spatial cognition and optimize reasoning in dynamic environments.
+Recently, SoFar (Qi et al., 2025) introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
 While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking.
-Inspired by prior spatial reasoning research [119, 114, 58, 76], we identified several limitations in existing benchmarks, such as reliance on generated images [104, 34, 48, 61, 92, 101], LLM-generated templates [59, 26, 91, 93], and domain-specific focus [117, 19, 21, 12, 28].
+Inspired by prior spatial reasoning research (Xu et al., 2024b, Wang et al., 2024a, Lin et al., 2014, Nwankwo et al., 2024), we identified several limitations in existing benchmarks, such as reliance on generated images (Szymanska et al., 2024, Fu et al., 2024b, Kamath et al., 2023, Liu et al., 2023a, Rajabi & Kosecka, 2023, Shiri et al., 2024), LLM-generated templates (Linghu et al., 2024, Du et al., 2024, Rädsch et al., 2025), and domain-specific focus (Xie et al., 2025, Chow et al., 2025, Danish et al., 2024).
 These issues hinder their comprehensiveness and real-world applicability.
@@
 6 Conclusion
-In this paper, we introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning.
-OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 1.5 K manually-curated question–answer pairs.
+We introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning.
+OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 8.4 K manually-curated question–answer pairs.
 Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking.
-By exposing these systematic gaps, OmniSpatial provides a rigorous yardstick and a rich error analysis playground for future work on physics- and viewpoint-aware multimodal models.
+To bridge these gaps, we introduce PointGraph for structured scene-graph reasoning and SpatialCoT for viewpoint-aware chain-of-thought, both yielding consistent gains and underscoring the value of structured and multi-view reasoning.
 Appendix A Detailed Task Design
@@
 B.1 Spatial Vision-Language Models
-Spatial vision-language models integrate computer vision [1, 37, 80, 90] and natural language processing [6, 115, 9, 2, 107, 108, 128] to enhance machine understanding of spatial relationships.
-Recent research [88, 15, 17, 11, 102, 123, 89] has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments.
-One of the pioneering works in this domain, SpatialVLM [15], has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset.
-These models [45, 45, 97, 115] effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
-Further advancements [62, 60, 94, 105, 120, 99, 116, 103] include SpatialRGPT [17], which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities.
-Similarly, SpatialBot [11] explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments.
-Recently, SoFar [89] proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities [18, 25, 73, 131, 113, 43, 74, 27, 42, 89].
+Spatial vision-language models integrate computer vision (He et al., 2022, Oquab et al., 2024, Radford et al., 2021) and natural language processing (Bai et al., 2023, Wang et al., 2024b, Brown et al., 2020, Anil et al., 2023, Touvron et al., 2023a; b, Zhang et al., 2024b) to enhance machine understanding of spatial relationships.
+Recent research (Qi et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Yuan et al., 2024, Qi et al., 2025) has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments.
+One of the pioneering works in this domain, SpatialVLM (Chen et al., 2024), has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset.
+These models (Hurst et al., 2024; 2024, Reid et al., 2024, Wang et al., 2024b) effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
+Further advancements (Liu et al., 2024b; a, Ramakrishnan et al., 2024, Tang et al., 2024, Yang et al., 2023, Rozanova et al., 2021, Wu et al., 2024, Stogiannidis et al., 2025) include SpatialRGPT (Cheng et al., 2024), which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities.
+Similarly, SpatialBot (Cai et al., 2024) explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments.
+Recently, SoFar (Qi et al., 2025) proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities (Cho et al., 2024, Driess et al., 2023, Mu et al., 2023, Zitkovich et al., 2023, Wake et al., 2024, Huang et al., 2024b, Nasiriany et al., 2024, Duan et al., 2024, Huang et al., 2024a, Qi et al., 2025, He et al., 2025a, Zhang et al., 2025b).
 However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks.
@@
 B.2 Spatial Reasoning in Psychology
-In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings [35].
+In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings (Gardner, 2011).
 To systematically characterize this ability, Buckley et al.
-[10]proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation.
+(2018)proposed a factor analysis-based framework, distinguishing between spatial visualization, spatial relations (such as mental rotation), and spatial orientation.
 Malanchini et al.
-[69] further identified a strong correlation between spatial orientation and object manipulation skills.
-Additionally, Newcombe & Shipley [75]introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic.
-Furthermore, many studies also have contributed to the development of frameworks [112, 52, 69, 39, 70] and evaluation methods [29, 111, 110, 40] for spatial reasoning.
+(2020) further identified a strong correlation between spatial orientation and object manipulation skills.
+Additionally, Newcombe & Shipley (2014)introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic.
+Furthermore, many studies also have contributed to the development of frameworks (Wai et al., 2009, Lee & Bednarz, 2012, Malanchini et al., 2020, Hegarty, 2010, McGarvey et al., 2018) and evaluation methods (Eliot & Smith, 1983, Uttal et al., 2024; 2013, Hegarty & Waller, 2005) for spatial reasoning.
 Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning.
@@
 Inspired by these classifications, our study proposes a novel taxonomy for visual-spatial reasoning, aiming to advance spatial reasoning research in vision-language models.
-Appendix C Additional Ablation Studies
-C.1 Evaluation and Format Strategy
-The choice of evaluation format and answer-extraction strategy can materially influence how fully a large language model’s capabilities are expressed.
-Because our benchmark introduces a novel visual–spatial reasoning task, Table 5 reports the performance impact of different prompting and evaluation schemes.
-We consider three prompting paradigms—Direct QA, Zero-shot CoT, and Manual CoT—and three evaluation methods: (i) regex-based field extraction (RE), in which the VLM is instructed to place its answer in a fixed slot at the end of the response; (ii) JSON extraction, where the answer must appear in a dedicated answer field; and (iii) LLM-based free-form evaluation, for which we employ GPT-4.1-mini as an adjudicator.
-We observe that reasoning-centric models that strongly obey formatting directives (e.g., Gemini-2.5-flash) cannot be reliably scored with regex extraction, because their compulsory chain-of-thought interferes with the required output template.
-For standard models, all three evaluation strategies yield comparable scores, and both CoT variants consistently outperform the Direct QA baseline by a small margin.
-Accordingly, in the main results table we evaluate non-reasoning models with the Manual CoT + RE setting, whereas reasoning-oriented models are assessed with Manual CoT + LLM evaluation.
-Promopt Type Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+Table 7: Comparative analysis of various prompting and evaluation strategies.
+Promopt Type
+Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
 Manipulate
@@
 Manual CoT LLM 36.26 53.78 34.10 50.35 42.10 44.00 23.51 26.06 60.39 31.38 23.86
-Table 5: Comparative analysis of various prompting and evaluation strategies.
+Appendix C Additional Ablation Studies
+C.1 Evaluation and Format Strategy
+The choice of evaluation format and answer-extraction strategy can materially influence how fully a large language model’s capabilities are expressed.
+Because our benchmark introduces a novel visual–spatial reasoning task, Table˜7 reports the performance impact of different prompting and evaluation schemes.
+We consider three prompting paradigms—Direct QA, Zero-shot CoT, and Manual CoT—and three evaluation methods: (i) regex-based field extraction (RE), in which the VLM is instructed to place its answer in a fixed slot at the end of the response; (ii) JSON extraction, where the answer must appear in a dedicated answer field; and (iii) LLM-based free-form evaluation, for which we employ GPT-4.1-mini as an adjudicator.
+We observe that reasoning-centric models that strongly obey formatting directives (e.g., Gemini-2.5-flash) cannot be reliably scored with regex extraction, because their compulsory chain-of-thought interferes with the required output template.
+For standard models, all three evaluation strategies yield comparable scores, and both CoT variants consistently outperform the Direct QA baseline by a small margin.
+Accordingly, in the main results table, we evaluate non-reasoning models with the Manual CoT + RE setting, whereas reasoning-oriented models are assessed with Manual CoT + LLM evaluation.
+C.2 Evaluation on the full OmniSpatial benchmark
+To complement the main experiments, we further conduct evaluation on the entire 8.4K OmniSpatial test set without applying the train/test split.
+This setting measures model performance on all available annotated questions, thereby reducing variance due to subset sampling and providing a more holistic view of spatial reasoning ability.
+Table 8 reports the results across all twelve fine-grained tracks.
+Overall, the relative ranking of models remains stable compared to the smaller-scale splits discussed in the main text.
+Gemini-2.5-Pro achieves the highest average score of 55.05, securing Rank 1 among all compared systems.
+ChatGPT o3 follows closely with 54.52, while Gemini-2.5-Flash achieves 51.80.
+Other proprietary systems, such as GPT-4.1 and GPT-4o, also maintain consistent performance.
+On the open-source side, Qwen-VL-2.5-32B achieves the strongest result with 43.97, outperforming other Qwen variants and aligning with the trends observed in the evaluations.
+A closer per-track analysis highlights complementary strengths across models.
+ChatGPT o3 excels at Manipulate, Traffic Analysis, and also leads in Pattern Recognition, Ego-Centric, and Hypothetical reasoning.
+In contrast, Gemini-2.5-Pro dominates Locate, Geospatial Strategy, Geometric Reasoning, and Allo-Centric perspective taking.
+Gemini-2.5-Flash yields the highest accuracy on Motion Analysis.
+These complementary strengths suggest that different architectures capture distinct aspects of spatial reasoning, and point towards potential benefits of ensembling or multi-expert distillation.
+Table 8: Evaluation on the full OmniSpatial benchmark. Dark green marks the best Avg. accuracy and light green marks the second-best Avg. accuracy in all the models.
+Method
+Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+Manipulate
+Motion
+Analysis
+Traffic
+Analysis
+Locate
+Geospatial
+Strategy
+Pattern
+Recognition
+Geometric
+Reasoning
+Ego
+Centric
+Allo
+Centric
+Hypothetical
+Proprietary Models
+GPT-4o 44.47 5 60.00 57.97 62.38 55.24 59.09 29.90 31.01 42.06 24.34 28.05
+GPT-4.1 48.75 4 61.18 66.67 66.60 69.52 60.91 30.93 35.58 47.06 27.26 33.06
+o3-04-16 54.52 2 71.05 63.64 70.61 69.05 61.70 42.55 47.15 53.35 30.98 46.28
+Gemini-2.0-Flash 42.41 7 56.47 56.76 55.95 63.81 57.27 15.46 30.04 43.20 24.56 28.05
+Gemini-2.5-Flash 51.80 3 54.84 70.43 65.33 66.00 68.75 39.53 39.46 52.70 33.49 36.84
+Gemini-2.5-Pro 55.05 1 66.67 68.34 69.10 78.85 71.15 36.17 50.17 50.65 39.57 37.89
+Open-source Models
+Qwen-VL-2.5-3B 39.76 8 57.65 47.34 49.07 49.52 58.18 29.90 26.97 40.62 30.47 29.88
+Qwen-VL-2.5-32B 43.97 6 58.82 52.42 52.97 68.57 51.82 27.84 29.51 51.70 26.82 35.98
+C.3 Details of Human Baseline & Inter-Annotator Agreement
+To establish an upper bound of performance, we conducted a human evaluation on the OmniSpatial benchmark.
+Six human annotators (graduate students with backgrounds in vision and robotics) were recruited.
+Each participant was presented with a randomized subset of questions covering all four tracks and a balanced selection of fine-grained tasks.
+The interface was blinded: multiple-choice options were randomized and no additional context was provided.
+Annotators were instructed to answer independently without external resources.
+Each item was labeled by three different annotators, and the final answer was obtained via majority vote.
+To quantify annotation reliability, we report inter-annotator agreement (IAA) using Krippendorff’s $\alpha$ for multi-annotator agreement.
+Table 9 summarizes the results across tracks.
+We also report 95% confidence intervals (bootstrap over questions) for human accuracy.
+Table 9: Human baseline accuracy and inter-annotator agreement (IAA) across tracks.
+Track Accuracy (%) Krippendorff’s $\alpha$
+Dynamic Reasoning 95.2 $\pm$ 1.3 0.92
+Spatial Interaction 93.5 $\pm$ 1.6 0.85
+Complex Logic 87.9 $\pm$ 5.5 0.76
+Perspective Taking 94.4 $\pm$ 2.2 0.80
+Overall 92.6 $\pm$ 1.5 0.84
+These results indicate that human annotators achieve $\sim$89% average accuracy, with substantial agreement across annotators ($\kappa$$\alpha\approx 0.84$).
+This provides a reliable estimate of the human upper bound and confirms the internal consistency of the benchmark.
 Appendix D System Prompts
-We present all the system prompts used in our experiments in Figs.
-7 and 8 to facilitate reproducibility.
+We present all the system prompts used in our experiments in Figs.˜7 and 8 to facilitate reproducibility.
 We observe that some models are sensitive to the choice of system prompt, which may stem from distributional biases in their training data.
@@
 Appendix E Additional Visualization
-We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs.
-9, 10, 11 and 12, respectively.
+We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs.˜9, 10, 11 and 12, respectively.
 Our benchmark comprises a rich collection of data samples spanning diverse scenarios, resolutions, lighting conditions, and geographical regions.
@@
 Appendix F Limitation & Future Works
-Although OmniSpatial includes some image clips with dynamic information from HOI4D [67], the complexity of the operational tasks still lags behind that of long videos [121, 14].
+Although OmniSpatial includes some image clips with dynamic information from HOI4D (Liu et al., 2022), the complexity of the operational tasks still lags behind that of long videos (Yang et al., 2024, Chandrasegaran et al., 2024).
 Moreover, while PointGraph & Spatial CoT enhances VLM’s spatial understanding through point cues, the improvement is not fundamental in nature.
-Spatial reasoning tasks are more like mathematics and coding tasks, require longer and more complex reasoning [22, 46].
-3D information is crucial for spatial reasoning, and future work involves introducing 3D representation [23, 86, 88, 87] and perception [84, 85, 127, 30], as well as 3D VLMs [81, 24, 36, 88, 89], reasoning model [125, 22, 79] and knowledge distillation [126, 41, 23, 86].
-The ultimate goal of spatial reasoning is to empower robots, and future work also involves robot execution tasks [42, 31, 89, 43, 38].
+Spatial reasoning tasks are more like mathematics and coding tasks, require longer and more complex reasoning (DeepSeek-AI et al., 2025, Jaech et al., 2024).
+3D information is crucial for spatial reasoning, and future work involves introducing 3D representation (Dong et al., 2023, Qi et al., 2023a; 2024; b) and perception (Qi et al., 2017a; b, Zhang et al., 2024a, Fan et al., 2024), as well as 3D VLMs (Peng et al., 2024, Dong et al., 2024, Guo et al., 2023, Qi et al., 2024; 2025), reasoning model (Zhang et al., 2025a, DeepSeek-AI et al., 2025, OpenAI, 2025b) and knowledge distillation (Zhang et al., 2023, Hinton et al., 2015, Dong et al., 2023, Qi et al., 2023a).
+The ultimate goal of spatial reasoning is to empower robots, and future work also involves robot execution tasks (Huang et al., 2024a, Fang et al., 2024a, Qi et al., 2025, Huang et al., 2024b, He et al., 2025b).
 Appendix G Broader Impacts

### v3

@@
-OmniSpatial: Towards Comprehensive Spatial Reasoning Benchmark for Vision Language Models
+arXiv:2506.03135v3[cs.CV] 28 Feb 2026
 Abstract
@@
 These models have achieved over 90% accuracy on previous benchmarks such as SpatialBot-Bench (Cai et al., 2024) and EmbSpatial (Du et al., 2024), suggesting that these basic tasks are approaching saturation.
-We believe complex spatial reasoning remains a significant challenge (Gardner, 2011, Baddeley, 1998, Previc, 1998, Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025).
-Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships (Bar-Anan et al., 2006, Trope & Liberman, 2010, Ramalho et al., 2018).
+We believe complex spatial reasoning remains a significant challenge (Gardner, 2011, Baddeley, 1998, Previc, 1998, Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025a).
+Human interaction with the physical world often involves interpreting ambiguous, dynamic, and context-dependent spatial relationships (Bar-Anan et al., 2006, Trope and Liberman, 2010, Ramalho et al., 2018).
 For example, in an emergency, knowing that an AED is “to the right of the door” is insufficient without understanding schematic diagrams, correlating maps with real-world environments, and planning an efficient route.
@@
 Existing benchmarks are nearly saturated on simple tasks like left–right discrimination, front–back identification, and object counting (Chen et al., 2024, Cai et al., 2024, Du et al., 2024, Szymanska et al., 2024).
-Yet real-world embodied tasks demand richer reasoning (Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025, Stogiannidis et al., 2025, Lee et al., 2025) about scene dynamics, multi-step logic, physical interaction, and viewpoint transformation.
+Yet real-world embodied tasks demand richer reasoning (Kosoy et al., 2025, Pothiraj et al., 2025, Chen et al., 2025a, Stogiannidis et al., 2025, Lee et al., 2025) about scene dynamics, multi-step logic, physical interaction, and viewpoint transformation.
 Guided by these considerations, we partition visual–spatial reasoning into four complementary dimensions: dynamic reasoning, complex spatial logic, spatial interaction, and perspective taking.
@@
 Together, these dimensions provide a framework for evaluating spatial reasoning in AI, robotics, and human cognition.
-Figure 2: Tasks Demonstration of OmniSpatial.
-Several representative subtasks are selected for demonstration in each of the four task categories.
-Note: The questions above are slightly simplified for clarity and conciseness.
+Figure 2: Tasks Demonstration of OmniSpatial. Several representative subtasks are selected for demonstration in each category. Note that the questions above are slightly simplified for clarity.
 3 OmniSpatial: Comprehensive Spatial Reasoning Benchmark
 3.1 Overview
-We present OmniSpatial, a comprehensive benchmark designed to evaluate vision-language models on spatial reasoning.
+We present OmniSpatial, a comprehensive benchmark designed to evaluate VLMs on spatial reasoning.
 Rather than pursuing sheer data volume, OmniSpatial emphasizes diversity, structure, and rigor.
-It now consists of 8.4K carefully curated question–answer pairs, substantially larger than earlier prototypes, and covers a broad spectrum of scenarios that demand reasoning beyond pattern recognition.
+It now consists of 8.4K carefully curated QA pairs, substantially larger than earlier prototypes, and covers a broad spectrum of scenarios that demand reasoning beyond pattern recognition.
 The dataset integrates heterogeneous sources—web imagery, standardized cognitive tests, driving-exam questions, and prior dataset images such as MME (Liang et al., 2024) and HOI4D (Liu et al., 2022).
@@
 Furthermore, to extend our dataset to model training, we partition the dataset into a 1.5K test set and a 6.9K training set.
-The test set is entirely human-annotated and further integrates samples from several existing datasets.
-including SpatialViz  (Wang et al., 2025), PhysBench (Chow et al., 2025), ViewSpatial (Li et al., 2025), and DrivingVQA (Corbière et al., 2025), which substantially enhance the diversity of our dataset.
+The test set is entirely human-annotated and the train set further integrates samples from several existing datasets, including SpatialViz  (Wang et al., 2025), PhysBench (Chow et al., 2025), ViewSpatial (Li et al., 2025), and DrivingVQA (Corbière et al., 2025), which substantially enhance the diversity.
 Figure 4: PointGraph: Enhance Spatial Reasoning through Additional Scene Graphs.
@@
 Gemini-2.5-Pro (Reid et al., 2024), InternVL3-78B (Zhu et al., 2025), and SoFar (Qi et al., 2025) achieve the best performance within their respective groups.
-Method
-Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+Method Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
 Manipulate
@@
 Random Choice 24.98 - 24.86 26.30 25.88 23.43 27.27 21.44 24.77 22.55 24.84 25.78
-GPT-3.5-turbo (Roumeliotis & Tselikas, 2023) 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
+GPT-3.5-turbo (Roumeliotis and Tselikas, 2023) 30.67 - 38.38 29.19 38.35 28.76 36.91 0.82 24.00 42.16 33.67 35.90
 GPT-4-turbo (OpenAI, 2023) 34.06 - 42.97 37.40 41.18 28.95 40.00 22.27 26.32 31.37 33.99 35.42
@@
 InternVL3-78B (Zhu et al., 2025) 49.33 1 63.78 63.12 56.24 59.24 51.45 27.63 30.19 74.51 38.46 35.90
-Qwen-VL2.5-3B (Wang et al., 2024b) 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
-Qwen-VL2.5-7B (Wang et al., 2024b) 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
-Qwen-VL2.5-32B (Wang et al., 2024b) 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
-Qwen-VL2.5-72B (Wang et al., 2024b) 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
+Qwen-VL2.5-3B (Wang et al., 2024c) 40.30 10 55.41 47.51 46.12 42.29 44.73 32.16 23.87 59.41 33.30 30.84
+Qwen-VL2.5-7B (Wang et al., 2024c) 39.18 12 58.38 35.09 50.12 45.33 44.00 31.13 29.42 64.51 33.19 37.35
+Qwen-VL2.5-32B (Wang et al., 2024c) 47.36 4 63.06 55.09 51.76 66.29 56.91 26.39 27.48 68.04 37.50 40.24
+Qwen-VL2.5-72B (Wang et al., 2024c) 47.85 3 58.38 60.12 50.12 59.81 53.64 26.19 33.03 71.37 36.81 36.39
 Specialized Spatial Reasoning Models
@@
 Table 3: Camparison of textual Chain-of-Thought and PointGraph on OmniSpatial-test.
-Method
-Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+Method Avg. Improve Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
 Manipulate
@@
 Overall Model Performance  As illustrated in Table˜2, we observe the following findings: (i) Proprietary reasoning models, such as ChatGPT o3 (OpenAI, 2025b) and Gemini-2.5-pro (Anil et al., 2023), achieve the highest performance, surpassing a  56% overall success rate; however, there remains a significant gap compared to human-level understanding, and they require a lot of inference time and tokens.
-(ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B (Zhu et al., 2025) and Qwen-VL2.5-72B (Wang et al., 2024b) achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp.
+(ii) Open-source models also demonstrate competitive results, with large-scale models like InternVL3-78B (Zhu et al., 2025) and Qwen-VL2.5-72B (Wang et al., 2024c) achieving comparable performance to GPT-4.1-mini and Gemini-2.0-flash-exp.
 (iii) Specialized Spatial Reasoning Models, due to limitations in dataset coverage and model capacity, struggle to achieve substantial improvements on comprehensive benchmarks.
@@
 Results in Table˜3 show a clear accuracy boost, particularly in the Dynamic Reasoning and Perspective-Taking Track, validating the benefits of integrating structured object representation, while traditional textual CoT difficult to bring about significant improvement.
-Fig.˜5 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track.
+Fig.˜5 and Table˜4 further demonstrates the effectiveness of our proposed Spatial CoT on the OmniSpatial Perspective-Taking track.
 Through novel view synthesis facilitated by InstantMesh, both GPT-4.1 and Qwen-VL2.5-7B exhibit significant performance improvements, validating the effectiveness of explicit spatial imagination.
@@
 RoboSpatial (Song et al., 2024) proposes a large-scale template-based spatial relationship dataset, focusing on positional relationships from different perspectives.
-VSI-Bench (Yang et al., 2024) combined video data (Chandrasegaran et al., 2024, Liu et al., 2023c, Fu et al., 2024a, Li et al., 2024b, Fang et al., 2024b) with cognitive maps (Momennejad et al., 2023, Apostolopoulos & Groumpos, 2023) to simulate human-like spatial cognition and optimize reasoning in dynamic environments.
+VSI-Bench (Yang et al., 2024) combined video data (Chandrasegaran et al., 2024, Liu et al., 2023c, Fu et al., 2024a, Li et al., 2024b, Fang et al., 2024b) with cognitive maps (Momennejad et al., 2023, Apostolopoulos and Groumpos, 2023) to simulate human-like spatial cognition and optimize reasoning in dynamic environments.
 Recently, SoFar (Qi et al., 2025) introduced the 6-DoF SpatialBench to evaluate the understanding of orientation.
 While these benchmarks have made significant contributions, a unified framework encompassing a wide range of complex spatial reasoning tasks remains lacking.
-Inspired by prior spatial reasoning research (Xu et al., 2024b, Wang et al., 2024a, Lin et al., 2014, Nwankwo et al., 2024), we identified several limitations in existing benchmarks, such as reliance on generated images (Szymanska et al., 2024, Fu et al., 2024b, Kamath et al., 2023, Liu et al., 2023a, Rajabi & Kosecka, 2023, Shiri et al., 2024), LLM-generated templates (Linghu et al., 2024, Du et al., 2024, Rädsch et al., 2025), and domain-specific focus (Xie et al., 2025, Chow et al., 2025, Danish et al., 2024).
+Inspired by prior spatial reasoning research (Xu et al., 2024b, Wang et al., 2024a, Lin et al., 2014, Nwankwo et al., 2024, Wang et al., 2024b), we identified several limitations in existing benchmarks, such as reliance on generated images (Szymanska et al., 2024, Fu et al., 2024b, Kamath et al., 2023, Liu et al., 2023a, Rajabi and Kosecka, 2023, Shiri et al., 2024), LLM-generated templates (Linghu et al., 2024, Du et al., 2024, Rädsch et al., 2025), and domain-specific focus (Xie et al., 2025, Chow et al., 2025, Danish et al., 2024).
 These issues hinder their comprehensiveness and real-world applicability.
 To address these gaps, our study proposes a comprehensive and integrative spatial reasoning benchmark.
+5.2 Spatial Vision-Language Models
+Spatial vision-language models integrate computer vision (He et al., 2022, Oquab et al., 2024, Radford et al., 2021) and natural language processing (Bai et al., 2023, Wang et al., 2024c, Brown et al., 2020, Anil et al., 2023, Touvron et al., 2023a; b, Zhang et al., 2024b) to enhance machine understanding of spatial relationships and mental intelligence.
+Recent research (Qi et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Yuan et al., 2024, Qi et al., 2025, Chen et al., 2025b) has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments.
+One of the pioneering works in this domain, SpatialVLM (Chen et al., 2024), has significantly advanced spatial reasoning by constructing an RGBD-based visual question-answering dataset.
+GS-Reasoner Chen et al.
+(2025b) introduces 3D grounding as a chain-of-thoughts for spatial reasoning tasks.
+These models (Hurst et al., 2024, Reid et al., 2024, Wang et al., 2024c) effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
+Further advancements (Liu et al., 2024b; a, Ramakrishnan et al., 2024, Tang et al., 2024, Yang et al., 2023, Rozanova et al., 2021, Wu et al., 2024, Stogiannidis et al., 2025) include SpatialRGPT (Cheng et al., 2024), which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities.
+Similarly, SpatialBot (Cai et al., 2024) explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments.
+Recently, SoFar (Qi et al., 2025) proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of vision-language models, significantly improving spatial understanding and robotic operation capabilities (Cho et al., 2024, Driess et al., 2023, Mu et al., 2023, Huang et al., 2024b, Nasiriany et al., 2024, Huang et al., 2024a, Qi et al., 2025, He et al., 2025a, Zhang et al., 2025b).
+However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks.
+To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
 6 Conclusion
 We introduce OmniSpatial, a benchmark for comprehensive visual–spatial reasoning.
-OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex spatial logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 8.4 K manually-curated question–answer pairs.
+OmniSpatial distills spatial cognition into four primary categories—dynamic reasoning, complex logic, spatial interaction, and perspective-taking—spanning 50 fine-grained subtasks and 8.4 K manually-curated question–answer pairs.
 Extensive experiments show that state-of-the-art proprietary and open-source VLMs peak at 57% accuracy—over 30 points below human performance—struggling especially with geometric reasoning and non-egocentric perspective taking.
-To bridge these gaps, we introduce PointGraph for structured scene-graph reasoning and SpatialCoT for viewpoint-aware chain-of-thought, both yielding consistent gains and underscoring the value of structured and multi-view reasoning.
+To bridge these gaps, we introduce PointGraph for structured scene-graph reasoning and SpatialCoT for viewpoint-aware CoT, both yielding consistent gains and underscoring the value of structured and multi-view reasoning.
+Ethic Statement
+This work complies with the ICLR Code of Ethics.
+All datasets are publicly available and used under their respective licenses.
+The research raises no direct ethical or legal concerns, and the authors are committed to responsible and fair use of the proposed methods.
+Reproducibility Statement
+We have made every effort to ensure the reproducibility of our work.
+The proposed model, implementation details, and evaluation protocols are described in detail in the main paper and appendix.
+All datasets used are publicly available and properly referenced.
+To further support reproducibility, we release the source code in the supplementary materials.
 Appendix A Detailed Task Design
@@
 Appendix B Additional Related Works
-B.1 Spatial Vision-Language Models
-Spatial vision-language models integrate computer vision (He et al., 2022, Oquab et al., 2024, Radford et al., 2021) and natural language processing (Bai et al., 2023, Wang et al., 2024b, Brown et al., 2020, Anil et al., 2023, Touvron et al., 2023a; b, Zhang et al., 2024b) to enhance machine understanding of spatial relationships.
-Recent research (Qi et al., 2024, Chen et al., 2024, Cheng et al., 2024, Cai et al., 2024, Song et al., 2024, Yuan et al., 2024, Qi et al., 2025) has increasingly focused on extending vision-language models to support spatial reasoning in dynamic and 3D environments.
-One of the pioneering works in this domain, SpatialVLM (Chen et al., 2024), has significantly advanced spatial reasoning by constructing an RGB-D based visual question-answering dataset.
-These models (Hurst et al., 2024; 2024, Reid et al., 2024, Wang et al., 2024b) effectively process multimodal data containing spatial information, serving as a bridge between visual perception and linguistic reasoning.
-Further advancements (Liu et al., 2024b; a, Ramakrishnan et al., 2024, Tang et al., 2024, Yang et al., 2023, Rozanova et al., 2021, Wu et al., 2024, Stogiannidis et al., 2025) include SpatialRGPT (Cheng et al., 2024), which extends RGB-D spatial understanding by incorporating 3D scene graphs and spatial data to improve inference capabilities.
-Similarly, SpatialBot (Cai et al., 2024) explores hierarchical deep reasoning mechanisms to handle depth and spatial structures in complex environments.
-Recently, SoFar (Qi et al., 2025) proposed semantic orientation and trained an open-world orientation model to enhance the orientation understanding of VLMs, significantly improving spatial understanding and robotic operation capabilities (Cho et al., 2024, Driess et al., 2023, Mu et al., 2023, Zitkovich et al., 2023, Wake et al., 2024, Huang et al., 2024b, Nasiriany et al., 2024, Duan et al., 2024, Huang et al., 2024a, Qi et al., 2025, He et al., 2025a, Zhang et al., 2025b).
-However, existing research remains limited in addressing the full complexity and comprehensiveness of spatial reasoning tasks.
-To bridge this gap, our study aims to develop a more comprehensive benchmark that rigorously evaluates spatial reasoning capabilities.
-B.2 Spatial Reasoning in Psychology
+B.1 Spatial Reasoning in Psychology
 In psychology, spatial reasoning refers to an individual’s ability to acquire, organize, utilize, and adapt spatial knowledge, recognized as one of the nine primary reasonings (Gardner, 2011).
@@
 (2020) further identified a strong correlation between spatial orientation and object manipulation skills.
-Additionally, Newcombe & Shipley (2014)introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic.
-Furthermore, many studies also have contributed to the development of frameworks (Wai et al., 2009, Lee & Bednarz, 2012, Malanchini et al., 2020, Hegarty, 2010, McGarvey et al., 2018) and evaluation methods (Eliot & Smith, 1983, Uttal et al., 2024; 2013, Hegarty & Waller, 2005) for spatial reasoning.
+Additionally, Newcombe and Shipley (2014)introduced a classification system for spatial thinking, dividing spatial reasoning into two dimensions: intrinsic-extrinsic and static-dynamic.
+Furthermore, many studies also have contributed to the development of frameworks (Wai et al., 2009, Lee and Bednarz, 2012, Malanchini et al., 2020, Hegarty, 2010, McGarvey et al., 2018) and evaluation methods (Eliot and Smith, 1983, Uttal et al., 2024; 2013, Hegarty and Waller, 2005) for spatial reasoning.
 Their framework highlights the distinction between object-centric spatial properties and external reference frames, offering valuable insights for applications such as navigation and path planning.
@@
 Table 7: Comparative analysis of various prompting and evaluation strategies.
-Promopt Type
-Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+Promopt Type Eval Type Avg. Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
 Manipulate
@@
 Table 8: Evaluation on the full OmniSpatial benchmark. Dark green marks the best Avg. accuracy and light green marks the second-best Avg. accuracy in all the models.
-Method
-Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+Method Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
 Manipulate
@@
 Perspective Taking 94.4 $\pm$ 2.2 0.80
-Overall 92.6 $\pm$ 1.5 0.84
+Overall 92.6 $\pm$ 2.5 0.84
 These results indicate that human annotators achieve $\sim$89% average accuracy, with substantial agreement across annotators ($\kappa$$\alpha\approx 0.84$).
+Even the most abstract and complex forms of spatial reasoning achieve a consistency score of 0.76.
 This provides a reliable estimate of the human upper bound and confirms the internal consistency of the benchmark.
 Appendix D System Prompts
-We present all the system prompts used in our experiments in Figs.˜7 and 8 to facilitate reproducibility.
+We present all the system prompts used in our experiments in Figs.˜12 and 13 to facilitate reproducibility.
 We observe that some models are sensitive to the choice of system prompt, which may stem from distributional biases in their training data.
 In contrast, inference-oriented models generally exhibit stronger generalization capabilities and are less reliant on carefully crafted prompts.
+We conduct extensive experiments and iterative tuning on the system prompts, with the ultimate goal of objectively and faithfully evaluating each model’s actual capability without being confounded by prompt-induced bias.
 Appendix E Additional Visualization
-We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs.˜9, 10, 11 and 12, respectively.
+We present more examples of question-answer pairs from the OmniSpatial, with perspective taking, spatial interaction, dynamic reasoning, and complex logic, shown in Figs.˜14, 15, 16 and 17, respectively.
 Our benchmark comprises a rich collection of data samples spanning diverse scenarios, resolutions, lighting conditions, and geographical regions.
 It includes both absolute numerical analyses and relative spatial relationships, aiming to comprehensively evaluate the spatial reasoning capabilities of vision-language models.
-Appendix F Limitation & Future Works
+To the best of our knowledge, our spatial intelligence benchmark is the most diverse and comprehensive to date, and is entirely human-annotated, enabling a faithful evaluation of models’ visual–spatial reasoning capabilities without the confounding effects of templated patterns.
+Appendix F Additional Experiments
+F.1 The Synergy of PointGraph and SpatialCoT
+Table 10: Performance of Spatial CoT on OmniSpatial Perspective-Taking track.
+Method Avg. Improve
+Ego
+Centric
+Allo
+Centric
+Hypothetical
+GPT-4.1-mini
+(w/ Zero‑shot CoT) 45.56 - 69.41 40.11 40.96
+(w/ Spatial CoT) 47.58 +2.02 69.43 42.37 44.34
+(w/ Spatial CoT & PointGraph) 48.70 +3.14 71.09 42.52 44.77
+Qwen‑VL2.5‑3B
+(w/ Zero‑shot CoT) 40.89 - 59.61 36.54 37.59
+(w/ Spatial CoT) 42.90 +2.01 60.80 39.25 37.44
+(w/ Spatial CoT & PointGraph) 43.75 +2.86 62.79 38.64 37.74
+In this section, we investigate the synergy between our two plug-and-play components, SpatialCoT and PointGraph.
+As shown in Table 10, across both GPT-4.1-mini and Qwen-VL2.5-3B, augmenting SpatialCoT with PointGraph yields an additional gain of roughly 1 percentage point in overall performance.
+F.2 Failure Case Analysis
+Figure 7 illustrates a Spatial Interaction task instance from OmniSpatial, together with failure cases of two state-of-the-art commercial models, Gemini-2.5-Pro and ChatGPT-o3.
+We observe that, despite their strong capabilities, these models still struggle to reason over complex 3D scenes and to perform orientation and path analysis under imagined egocentric poses.
+In this example, the model must first imagine itself standing in front of the washing machine and then plan a path in the 3D scene graph accordingly.
+We hypothesize that, if the model could generate intermediate visualizations of its planned route (e.g., hand-drawn trajectories) as a form of chain-of-thought (CoT), mimicking human problem-solving, it would be more effective at handling such challenging cases.
+Figure 7: Qualitative example for Failure Studies.
+In Fig.
+8, we present a representative Perspective-Taking instance from OmniSpatial.
+We observe that even these top-performing models struggle with frame-of-reference confusion, counterfactual viewpoints and the associated spatial relations.
+In this example, the model must imagine itself facing the vase and then infer the relative positions of objects under this imagined viewpoint.
+Because current models find it difficult to construct and manipulate such fictitious perspectives, we propose using novel-view synthesis as a plug-and-play SpatialCoT module to enhance their Perspective-Taking ability.
+In future work, more tightly integrating this capability into the VLM’s internal thinking process may further strengthen the model’s spatial understanding.
+Figure 8: Qualitative example for Failure Studies.
+In Fig.
+9, we present an example spatio-temporal dynamics reasoning task from OmniSpatial.
+We observe that even state-of-the-art commercial models struggle to understand long-horizon sequences involving multiple continuous actions.
+We hypothesize that performance on such long-term temporal understanding tasks can be improved by augmenting the model’s memory, for example by enriching its temporal and episodic representations of time–space relations.
+Figure 9: Qualitative example for Failure Studies.
+In Figure 10, we present an example of a Complex-Logic task from OmniSpatial.
+In this problem, the model must understand the folding relationships in a paper-cutting process, which is highly challenging even for humans.
+We observe that the reasoning model produces a detailed chain of analysis, including examining how each folded corner affects the boundary and attempting to reason using structured text.
+However, we argue that such complex logical reasoning is better addressed with a vision-grounded chain-of-thought (CoT), i.e., a unified generative–perceptual model whose generative capability is used to predict the intermediate folded states step by step, gradually approaching the final correct prediction.
+Figure 10: Qualitative example for Failure Studies.
+In addition, we further conduct a quantitative analysis of how depth and object orientation affect the model’s accuracy.
+We collect all real-image samples from the viewpoint conversion task and analyze them using Gemini-Flash-2.5 Reid et al.
+(2024).
+Object depth is estimated with Metric3D v2 Yin et al.
+(2023) and discretized into several ranges, while object orientation is evaluated with SoFar Qi et al.
+(2025) and grouped by angle.
+As shown in Fig.
+11, we observe that the model’s accuracy slightly decreases as the distance between the object and the camera increases.
+Regarding orientation, the model handles oblique views better, whereas its understanding of rear-view objects is somewhat weaker.
+Figure 11: Quantitative example for Failure Studies.
+F.3 Main Result with Standard Deviation
+In Table 11, we additionally report results with standard deviations.
+Due to space limitations, we only present the means and standard deviations of the four coarse-grained categories, keeping the configuration consistent with the main table in the paper.
+We observe that the reasoning models exhibit larger variability across the five runs, whereas the smaller-parameter open-source models produce more consistent results.
+Table 11: Evaluation on OmniSpatial-test.
+All models were tested 5 times and averaged to reduce randomness.
+We report accuracy on four coarse-grained categories: Dynamic Reasoning, Spatial Interaction, Complex Logic, and Perspective Taking.
+Each coarse category is the arithmetic mean of its corresponding fine-grained sub-tasks.
+Method Avg. Rank Dynamic Reasoning Spatial Interaction Complex Logic Perspective Taking
+Proprietary Models
+GPT-4o-mini-2024-07-18 (Hurst et al., 2024) 42.64 $\pm$ 0.9 8 53.12 $\pm$ 1.1 47.64 $\pm$ 1.4 25.95 $\pm$ 0.4 44.18 $\pm$ 1.6
+GPT-4o-2024-11-20 (Hurst et al., 2024) 47.81 $\pm$ 1.0 5 61.39 $\pm$ 1.3 54.31 $\pm$ 1.6 25.88 $\pm$ 0.5 51.74 $\pm$ 1.8
+GPT-4.1-nano-2025-04-14 (OpenAI, 2025a) 42.62 $\pm$ 0.8 9 52.38 $\pm$ 1.0 46.09 $\pm$ 1.2 27.25 $\pm$ 0.4 41.52 $\pm$ 1.4
+GPT-4.1-mini-2025-04-14 (OpenAI, 2025a) 48.87 $\pm$ 1.0 4 60.42 $\pm$ 1.2 58.54 $\pm$ 1.5 29.73 $\pm$ 0.5 50.47 $\pm$ 1.9
+GPT-4.1-2025-04-14 (OpenAI, 2025a) 51.78 $\pm$ 1.1 2 65.48 $\pm$ 1.4 61.84 $\pm$ 1.7 30.91 $\pm$ 0.6 50.22 $\pm$ 1.7
+Claude-3-5-sonnet-20241022 (Anthropic, 2024) 46.86 $\pm$ 0.9 6 54.31 $\pm$ 1.1 59.86 $\pm$ 1.3 29.17 $\pm$ 0.4 48.10 $\pm$ 1.5
+Claude-3-7-sonnet-20250219 (Anthropic, 2024) 47.53 $\pm$ 0.9 5 56.76 $\pm$ 1.2 59.87 $\pm$ 1.4 28.94 $\pm$ 0.5 48.28 $\pm$ 1.6
+Gemini-2.0-flash-lite-02-05 (Anil et al., 2023) 44.03 $\pm$ 0.8 8 52.95 $\pm$ 1.0 54.34 $\pm$ 1.2 26.44 $\pm$ 0.4 47.36 $\pm$ 1.4
+Gemini-2.0-flash-exp (Anil et al., 2023) 48.40 $\pm$ 1.0 2 58.95 $\pm$ 1.3 58.09 $\pm$ 1.6 27.32 $\pm$ 0.5 50.41 $\pm$ 1.8
+Gemini-2.5-flash-preview-05-20 (Anil et al., 2023) 52.12 $\pm$ 1.1 1 65.14 $\pm$ 1.5 67.49 $\pm$ 1.7 36.16 $\pm$ 0.6 48.37 $\pm$ 1.9
+Reasoning Models
+o1-2024-12-17 (Jaech et al., 2024) 50.36 $\pm$ 1.1 6 66.30 $\pm$ 1.1 60.49 $\pm$ 1.3 33.14 $\pm$ 1.3 48.58 $\pm$ 1.7
+o4-mini-2025-04-16 (OpenAI, 2025b) 52.77 $\pm$ 1.2 3 66.40 $\pm$ 1.3 65.05 $\pm$ 1.5 35.40 $\pm$ 1.5 51.73 $\pm$ 1.9
+o3-2025-04-16 (OpenAI, 2025b) 56.33 $\pm$ 1.3 1 69.03 $\pm$ 1.4 65.07 $\pm$ 1.6 34.95 $\pm$ 1.7 57.88 $\pm$ 2.0
+Claude-3-7-sonnet-20250219-thinking (Anthropic, 2024) 48.62 $\pm$ 1.0 7 58.47 $\pm$ 1.0 59.65 $\pm$ 1.2 29.20 $\pm$ 1.1 47.84 $\pm$ 1.5
+Gemini-2.5-flash-05-20-thinking (Anil et al., 2023) 53.16 $\pm$ 1.2 3 67.50 $\pm$ 1.2 63.91 $\pm$ 1.4 35.59 $\pm$ 1.4 49.20 $\pm$ 1.8
+Gemini-2.5-pro-preview-05-06 (Anil et al., 2023) 55.19 $\pm$ 1.4 2 69.48 $\pm$ 1.5 67.38 $\pm$ 1.7 39.07 $\pm$ 1.8 49.96 $\pm$ 2.0
+Open-source Models
+LLavA-1.5-vicuna-7B (Liu et al., 2024c) 34.97 $\pm$ 0.7 15 42.84 $\pm$ 0.9 35.14 $\pm$ 1.1 26.59 $\pm$ 0.4 42.13 $\pm$ 1.3
+LLaVA-onevision-qwen2-7B (Li et al., 2024a) 35.68 $\pm$ 0.7 14 40.70 $\pm$ 0.8 34.76 $\pm$ 1.0 25.73 $\pm$ 0.3 40.19 $\pm$ 1.2
+LLaVA-onevision-qwen2-72B (Li et al., 2024a) 45.66 $\pm$ 0.9 6 56.22 $\pm$ 1.1 57.14 $\pm$ 1.3 24.24 $\pm$ 0.5 49.14 $\pm$ 1.6
+Gemma-3-4B (Kamath et al., 2025) 39.79 $\pm$ 0.8 11 45.80 $\pm$ 0.9 40.15 $\pm$ 1.2 24.12 $\pm$ 0.4 44.84 $\pm$ 1.4
+Gemma-3-12B (Kamath et al., 2025) 43.71 $\pm$ 0.8 8 54.48 $\pm$ 1.0 49.06 $\pm$ 1.3 23.41 $\pm$ 0.4 44.72 $\pm$ 1.5
+Gemma-3-27B (Kamath et al., 2025) 44.75 $\pm$ 0.9 7 56.27 $\pm$ 1.1 53.62 $\pm$ 1.4 28.44 $\pm$ 0.5 43.58 $\pm$ 1.6
+InternVL3-2B (Zhu et al., 2025) 37.98 $\pm$ 0.7 13 45.29 $\pm$ 0.8 41.28 $\pm$ 1.1 25.19 $\pm$ 0.3 41.20 $\pm$ 1.3
+InternVL3-8B (Zhu et al., 2025) 41.60 $\pm$ 0.8 9 46.65 $\pm$ 0.9 48.25 $\pm$ 1.2 26.79 $\pm$ 0.4 47.93 $\pm$ 1.4
+InternVL3-14B (Zhu et al., 2025) 45.94 $\pm$ 0.8 5 57.25 $\pm$ 1.0 51.20 $\pm$ 1.3 28.15 $\pm$ 0.4 45.96 $\pm$ 1.5
+InternVL3-38B (Zhu et al., 2025) 48.48 $\pm$ 1.0 2 63.50 $\pm$ 1.2 54.48 $\pm$ 1.4 29.21 $\pm$ 0.5 47.47 $\pm$ 1.7
+InternVL3-78B (Zhu et al., 2025) 49.33 $\pm$ 1.0 1 63.45 $\pm$ 1.3 55.64 $\pm$ 1.5 28.91 $\pm$ 0.5 49.62 $\pm$ 1.8
+Qwen-VL2.5-3B (Wang et al., 2024c) 40.30 $\pm$ 0.8 10 51.46 $\pm$ 1.0 44.38 $\pm$ 1.2 28.02 $\pm$ 0.4 41.18 $\pm$ 1.4
+Qwen-VL2.5-7B (Wang et al., 2024c) 39.18 $\pm$ 0.9 12 46.73 $\pm$ 1.0 46.48 $\pm$ 1.3 30.27 $\pm$ 0.6 45.02 $\pm$ 1.5
+Qwen-VL2.5-32B (Wang et al., 2024c) 47.36 $\pm$ 1.0 4 59.08 $\pm$ 1.2 58.32 $\pm$ 1.5 26.94 $\pm$ 0.5 48.59 $\pm$ 1.7
+Qwen-VL2.5-72B (Wang et al., 2024c) 47.85 $\pm$ 0.9 3 59.25 $\pm$ 1.2 54.52 $\pm$ 1.4 29.61 $\pm$ 0.5 48.19 $\pm$ 1.6
+Specialized Spatial Reasoning Models
+SpaceMantis-13B (Chen et al., 2024) 36.36 $\pm$ 0.7 6 41.81 $\pm$ 0.9 36.30 $\pm$ 1.1 23.33 $\pm$ 0.4 42.25 $\pm$ 1.3
+SpaceQwen2.5-VL-3B (Chen et al., 2024) 40.25 $\pm$ 0.9 3 49.00 $\pm$ 1.0 41.01 $\pm$ 1.2 27.85 $\pm$ 0.5 47.44 $\pm$ 1.6
+SpaceThinker-Qwen2.5VL-3B (Chen et al., 2024) 40.42 $\pm$ 0.9 2 50.45 $\pm$ 1.1 39.15 $\pm$ 1.3 26.16 $\pm$ 0.5 41.41 $\pm$ 1.4
+SpatialBot-3B (Cai et al., 2024) 35.68 $\pm$ 0.7 6 40.70 $\pm$ 0.8 34.76 $\pm$ 1.0 25.73 $\pm$ 0.3 40.19 $\pm$ 1.2
+RoboPoint-vicuna-v1.5-7B-lora (Yuan et al., 2024) 35.85 $\pm$ 0.7 6 42.82 $\pm$ 0.9 37.57 $\pm$ 1.1 26.30 $\pm$ 0.4 43.29 $\pm$ 1.3
+RoboPoint-vicuna-v1.5-13B (Yuan et al., 2024) 34.60 $\pm$ 0.7 5 41.91 $\pm$ 0.9 35.85 $\pm$ 1.1 25.93 $\pm$ 0.4 40.06 $\pm$ 1.3
+SoFar-Qwen2.5VL-3B (Qi et al., 2025) 45.14 $\pm$ 1.0 1 53.83 $\pm$ 1.2 53.33 $\pm$ 1.4 27.31 $\pm$ 0.6 49.95 $\pm$ 1.7
+Human Evaluation
+Human 92.63 $\pm$ 2.5 - 95.24 $\pm$ 1.3 93.48 $\pm$ 1.6 87.86 $\pm$ 5.5 94.36 $\pm$ 2.2
+F.4 SpatialCoT on other track
+Besides the perspective-taking tasks, SpatialCoT also improves performance on many problems that require viewpoint transformation.
+We further evaluate SpatialCoT on the Complex-Logic track; as shown in Table 12, SpatialCoT consistently achieves better performance than Manual CoT.
+Table 12: Performance of Spatial CoT on Complex-Logic track.
+Method Avg. Improve
+Pattern
+Recognition
+Geometric
+Reasoning
+GPT-2.5-Flash-Thinking
+(w/ Manual CoT) 35.59 - 35.05 36.13
+(w/ Spatial CoT) 37.19 +1.60 35.30 38.28
+Qwen‑VL2.5‑3B
+(w/ Manual CoT) 28.02 - 32.16 23.87
+(w/ Spatial CoT) 29.27 +1.25 29.01 28.52
+Appendix G Limitation & Future Works
 Although OmniSpatial includes some image clips with dynamic information from HOI4D (Liu et al., 2022), the complexity of the operational tasks still lags behind that of long videos (Yang et al., 2024, Chandrasegaran et al., 2024).
@@
 The ultimate goal of spatial reasoning is to empower robots, and future work also involves robot execution tasks (Huang et al., 2024a, Fang et al., 2024a, Qi et al., 2025, Huang et al., 2024b, He et al., 2025b).
-Appendix G Broader Impacts
+Appendix H Broader Impacts
 OmniSpatial promises several positive societal benefits.